### v1

# Quantifying the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data

Abstract
Robustness has become a critical attribute for the deployment of RAG systems in real-world applications. Existing research focuses on robustness to explicit noise (e.g., document semantics) but overlooks spurious features (a.k.a. implicit noise). While previous works have explored spurious features in LLMs, they are limited to specific features (e.g., formats) and narrow scenarios (e.g., ICL). In this work, we statistically confirm the presence of spurious features in the RAG paradigm, a robustness problem caused by the sensitivity of LLMs to semantic-agnostic features. Moreover, we provide a comprehensive taxonomy of spurious features and empirically quantify their impact through controlled experiments. Further analysis reveals that not all spurious features are harmful and they can even be beneficial sometimes. Extensive evaluation results across multiple LLMs suggest that spurious features are a widespread and challenging problem in the field of RAG. We release all codes and data at: https://github.com/maybenotime/RAG-SpuriousFeatures.
Machine Learning, ICML
1 Introduction
Figure 1: An example from the SURE framework (Sec. 3), illustrating the sensitivity of RAG systems to spurious features within grounding data. MD represents the Markdown format. The original retrieved document is fed into the LLMs in different formats, leading to inconsistent responses.
Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm to mitigate LLMs hallucinations (Gao et al., 2023; Yang et al., 2023b), integrating relevant external knowledge to improve the factuality and trustworthiness of LLM-generated outputs (Zhou et al., 2024). However, Retrieval-Augmented Language Models (RALMs) still face substantial robustness issue due to the presence of noise in retrieved documents (Liu et al., 2023; Li et al., 2024c).
Recent research aims to explore the characteristics that affect the robustness of RAG systems from the perspective of grounding data construction (Cuconasu et al., 2024). These studies examine various factors, including the type (Wu et al., 2024a), number (Xu et al., 2024), and position of documents (Liu et al., 2024) within the prompt context. A more detailed discussion of related work is in Appendix A. However, previous analyses primarily focus on explicit noise (i.e., causal features) that significantly alter the semantic information of grounding data(Wu et al., 2024b; Cuconasu et al., 2024), while neglecting spurious features (a.k.a. implicit noise) that make slight modifications but still preserve the meaning of contexts. This limitation extends to existing evaluation benchmarks, which simulate complex noise scenarios to assess the robustness of RAG systems (Chen et al., 2024a; Wu et al., 2024a), yet lack available benchmarks and metrics to measure the robustness of LLMs against spurious features.
Contemporary RAG systems typically employ production-level retrievers, such as Bing and Google, to collect relevant information from the internet. Unlike a single corpus, the internet encompasses diverse data with distinct features. For any given query, there may exist numerous golden documents that contain the correct answer but differ in style, format, or other attributes. As shown in Figure 1, we have observed that LLMs may fail to consistently derive the correct answer from golden documents with different formats. A similar phenomenon is reported in  Sclar et al. (2024) and  He et al. (2024), which demonstrate that LLMs are extremely sensitive to the format of prompts (i.e., spurious features). For more related work, see Appendix A. Unfortunately, there is no statistic and empirical evidence to support the existence of spurious features in the RAG paradigm. This highlights the urgent need to redefine spurious features in RAG and systematically quantify the robustness of RALMs against them.
To address these challenges, we first design a preliminary experiment to demonstrate that RALMs are sensitive to semantic-agnostic features in the grounding data, thereby extending the definition of spurious features to RAG systems. Building on findings from our preliminary experiment and recent studies, we identify five types of spurious features that may appear in RAG scenarios. Then, we propose a novel framework, SURE, for automating the process of robustness evaluation. This framework follows a perturb-then-evaluate approach, offering great scalability. In SURE, automated perturbations are applied on the original instances to inject the corresponding spurious features. The perturbed instances are then examined to ensure that the causal features remain intact. After these steps, we employ tailored metrics to quantify the robustness of RALMs against spurious features. Further analysis reveals that not every spurious features is harmful and they can even be beneficial sometimes. We also find that our proposed metrics are more suitable for evaluating spurious features in golden documents. Based on this finding, we distill the most challenging instances from the synthetic data generated by our framework to create a lighter benchmark, SIG, enabling more efficient robustness evaluation. We evaluate 12 representative LLMs with varying architectures and scales, and the results show that maintaining robustness against spurious features remains a significant challenge, especially for open-source LLMs.
Our contribution can be summarized as follows: 1) We extend the definition of spurious features to RAG systems through a preliminary experiment. To the best of our knowledge, this is the first comprehensive study to define and evaluate spurious features from RAG perspective. 2) We propose a novel evaluation framework, SURE, to assess the robustness of RALMs against spurious features, which effectively simulate five types of spurious features that may appear in real world scenarios. 3) Through extensive experiments and analysis using SURE on two representative LLMs, we provide valuable insights for future research. 4) Based on these findings, we curate a lightweight yet challenging evaluation dataset named SIG and benchmark the robustness of the current state-of-the-art LLMs against spurious features.
2 Preliminary
In this section, we first define causal and spurious features in the context of retrieval-augmented generation and then provide statistical evidence to support the existence of spurious features. Specifically, we start by defining an oracle retriever to sample data according to the preferences of LLMs and then apply statistical testing to validate that RALMs exhibit biases toward semantic-agnostic features within the grounding data.
2.1 Causal and Spurious Features in RAG
In general, causal features are input features that have a direct causal effect on the output of predictive model (Yu et al., 2020). Their relationship is rooted in causality, rather than mere statistical correlation. When it comes to Large Language Models, the meaning and intent of prompts serve as causal features that directly influence the models’ responses. In the context of RAG, causal features refer to the semantic information of grounding data, including whether the correct answer exists, the amount of noise documents in the grounding data, and the content of those noise documents.
In contrast, spurious features are input features that co-occur with causal features and are erroneously captured by the model (Neuhaus et al., 2023). These features exhibit a statistical correlation with the model’s output but lack a causal relationship. Recent research has shown that LLMs are sensitive to seemingly trivial features like prompt formatting, thereby extending the definition of spurious features to LLMs (Sclar et al., 2024). Similarly, we hypothesize that the semantic-agnostic features of the grounding data can be defined as spurious features in RAG systems. However, the conclusion drawn from LLMs may not be applicable to RALMs. Unlike LLMs, the input prompts for RALMs incorporate additional dynamic content—grounding data—to augment their output. This grounding data is derived from retrieved documents rather than static instructions provided directly by users. Therefore, we design a preliminary experiment to validate whether RALMs are sensitive to semantic-agnostic features within the grounding data.
2.2 Oracle Retriever
We aim to confirm the existence of spurious features in RAG scenarios, i.e., to demonstrate the sensitivity of RALMs to semantic-agnostic features in grounding data.
There are some challenges in revealing the sensitivity of RALMs. Specifically, When retrieving from a single corpus, it is difficult to mine semantically equivalent counterparts with obvious differences in semantic-agnostic features. This is because documents from the corpus often share similar styles and formats. With only slight differences in these features, we are difficult to observe significant performance variations using traditional evaluation metrics like accuracy. Thus, more fine-grained metrics are required to capture the subtle performance changes. Inspired by the use of LLMs as supervision signals for document utility (Izacard et al., 2023; Gan et al., 2024), we propose the oracle score, which measures fine-grained performance through calculating the log probability of generating correct answers based on the given documents. The oracle score is defined as follows:
[…]
where $x$ is the input prompt for RALMs, including the instruction $I$, grounding data $G$, and query $Q$; $y$ represents the ground truth answer; $\theta$ denotes the model parameters; and $T$ is the total length of the answer sequence. For cases with multiple answers, we compute the final score by averaging the corresponding oracle scores across all answers. Notably, the oracle score is less effective at estimating the helpfulness of documents when handling queries with long-sequence answers  [^1] In our experiments, we only use queries where the answer sequences is fewer than 5 tokens, ensuring the effectiveness of oracle scores.. This limitation arises from the growing expansion of the solution space as the answer length increases.
We further define an oracle retriever that ranks documents according to their oracle scores. Since the oracle score is computed using LLMs as supervision signals, the oracle retriever can be regarded as a sampler that reflects the LLMs’ preferences for grounding data. With this sampler, we successfully transform the target of demonstrating that RALMs are sensitive to semantic-agnostic features within grounding data into showing that the oracle retriever is biased toward spurious features in documents.
2.3 Preliminary Experiment & Analysis
Using the oracle retriever, we recall 100 documents from the Wikipedia dump for each query in the NQ-open dataset. However, the computational cost of calculating oracle scores across the entire corpus to select top-ranked documents is prohibitive. Therefore, we introduce Contriever-msmarco, a traditional dense retriever, for first-stage retrieval, followed by reranking in descending order based on the oracle scores. In addition to reducing the computational load, this initial retrieval also ensures the semantic similarity of retrieved documents. To further eliminate the effect of causal features, documents without golden answers are filtered out, ensuring that the remaining documents have roughly consistent causal features.
We then select the first-ranked and last-ranked documents for each query from the remaining documents, resulting in two sets, each containing 2658 samples. By comparing the differences in feature distributions between these two sets, we can assess whether the oracle retriever exhibits bias toward semantic-agnostic features. If these two sets do not belong to the same feature distribution, this can be attributed to the oracle retriever’s bias towards semantic-agnostic features during sampling. To confirm that this bias is not introduced by the dense retriever in first-stage retrieval, we establish a control group by randomly sampling two documents instead of selecting the first- and last-ranked documents.
To evaluate whether the two distributions are same, we employ the Kolmogorov-Smirnov (K-S) test. The following semantic-agnostic features are measured in the experiments: 1) Flesh Score, 2) Distinct-1, 3) Dependency Tree Depth, 4) PPL, and 5) Token Length. A detailed introduction of the K-S test and these features can be found in the Appendix B.
We conduct experiments using Mistral-7B-Instruct-v0.3 to implement the oracle retriever. The K-S statistic and P-value are presented in Table 3. Furthermore, we visualize the feature distributions for both the experimental and control groups in Figure 6. For all tested features in the experimental group, the K-S test rejects the null hypothesis, concluding that the distribution of the two sets are significantly different. In contrast, for the control group, the K-S test fails to reject the null hypothesis. The results for Llama-3.1-8B-Instruct are also provided in Appendix B. According to these results, we can conclude that RALMs exhibit bias toward spurious features in documents.
The preliminary experiment provides statistical evidence supporting the existence of spurious features in RAG systems. Nevertheless, it does not offer empirical evidence or quantitative analysis. Inspired by previous data synthesis studies (Tan et al., 2024b; Tong et al., 2024; Li et al., 2024b; Wang et al., 2024a), we use a data synthesis approach to better control feature variables and quantify the robustness of RALMs against spurious features.
3 Proposed Framework
Figure 2: Overview of our SURE framework. We provide a Comprehensive Taxonomy that includes five types of spurious features, further divided into 13 subtypes of perturbations (left section). To construct the testbed, we prepare raw instances initially and then synthesize the modified instances through a workflow consisting of Spurious Features Injection and Causal Features Preservation (middle section). By applying carefully tailored metrics for Robustness Evaluation, we quantify the robustness of target RAG systems (right section).
In this section, we detail our proposed evaluation framework, SURE (Spurious FeatUres Robustness Evaluation), which designed specifically for assessing the robustness of RALMs against spurious features in grounding data. As illustrated in Figure 2, this framework comprise four components: 1) Comprehensive Taxonomy. We identify five distinct types of spurious features that may arise in the context of Retrieval Augmented Generation. 2) Spurious Features Injection. We design a data synthesis pipeline to automate the injection of spurious features, utilizing both model-based and rule-based methods to construct counterparts of the original document with varying spurious features. 3) Causal Features Preservation. We employ a bidirectional entailment algorithm and a string matching strategy to ensure that the causal features of grounding data remain unchanged. 4) Robustness Evaluation. We report the win, lose and robustness rate at instance level to facilitate a fine-grained evaluation.
3.1 Problem Formulation
Given a query $q$ , the retriever $R$ returns a list of relevant documents from a corpus $D=\{d_{i}\}_{i=1}^{N}$. The relevance between document $d$ and query $q$ can be measured by various methods. In this work, we use a BERT-based dense retriever to obtain the embedding of query and documents, respectively. The relevance score is calculated by computing their dot-product similarity:
$s(q,d_{i})=E(q)\,\raisebox{-0.86108pt}{$\cdot$}\,E(d_{i}).$ (2)
Then, the Top-k documents with the highest similarity scores are retrieved:
$D_{\text{retrieve}}=\text{argtop-}k\left\{s(q,d_{i})\mid d_{i}\in D\right\}.$ (3)
To formally quantify the robustness of RAG systems against spurious features, we define the input prompt for the LLM-based reader as $P=(I,G,Q)$, where $I$ represents instruction, $G$ refers to the grounding data, constituted by a subset of $D_{\text{retrieve}}$, and $Q$ is the query. A perturbation is introduced to investigate the impact of spurious features by applying a semantic-agnostic modification to the original grounding data, while preserving its causal features. We define $g(.)$ to automate this process, transforming $G$ to $g(G)$ and producing a counterpart $\hat{P}=(I,g(G),Q)$. The outputs of LLM-based reader for $P$ and $\hat{P}$ are compared to evaluate the impact of the introduced perturbation:
$y=\text{LLM}(P),\quad\hat{y}=\text{LLM}(\hat{P}).$ (4)
3.2 Taxonomy of Spurious Features
Based on the biased spurious features identified in our preliminary experiment and prior works, we introduce five types of spurious features and their corresponding perturbations in detail as follows.
Style Perturbations
The same content can be expressed in different styles, using varying tones, words and sentence structures. As shown in Section 2.3, LLMs exhibit biases towards readability-related features. Similarly, for humans, the readability of a text can significantly influence its accessibility to the audience (Yang et al., 2023a). Therefore, we define two perturbations from the perspective of readability style: Simple and Complex. The former simplifies the grounding data by using basic vocabulary and simple sentence structure, while the latter employs professional vocabulary and a formal academic tone to complex the documents.
Source Perturbations
LLM-generated content, including both misinformation and correct claims, infiltrates every corner of the internet. Recent studies have shown that neural retrievers are biased towards LLM-generated content, leading to the marginalization of human-authored content (Dai et al., 2024; Chen et al., 2024b). Moreover, our preliminary experiments demonstrate that LLMs are biased towards the Perplexity (PPL) of text. Thus, we define two types of source perturbations: LLM-generated and Self-generated. Specifically, the LLM-generated perturbation paraphrases the original document using a powerful LLM, while the self-generated perturbation employs the same backbone model used as the generator in the RAG system.
Logic Perturbations
The arrangement of sentences within a passage typically follows a logical order, ensuring the clarity and coherence in the narrative flow. Here, we simulate scenarios where the intrinsic logical chain is disrupted by three different perturbations: Random, Reverse, and LLM-reranked, each representing a distinct sentence ordering strategy.
Format Perturbations
The internet contains various data formats, including HTML, Markdown, YAML and JSON. These formats are usually processed into plain text before being fed to LLMs. To mitigate the loss of structural information during this process, some RAG studies propose using the original format, rather than plain text, to augment the generation (Tan et al., 2024a). However, as highlighted in previous research, the prompt format is recognized as a spurious feature that can significantly impact model performance (Sclar et al., 2024; He et al., 2024). Therefore, we perturb the original document with four common formats to explore the impact of grounding data format in the context of RAG.
Metadata Perturbations
Metadata is often included in the HTML results returned by search engines. In our framework, we consider two main types of metadata: Timestamp and Data source. The timestamp indicates the time when the data was created, and data source identifies the origin of the data.
3.3 Spurious Features Injection
The automation of spurious features injection is essential for automating the entire evaluation framework. We detail the process of collecting the original instances and describe how the automated perturbation was implemented.
Instance Preparation
An instance is the dynamic component of the prompt $P$, consisting of a query $Q$ and grounding data $G$. The queries are drawn from the NQ-open dataset, while our data source is English Wikipedia dump as of 20 December 2018. To construct the original instances, we first select 1,000 queries based on the close-book QA results of Mistral-7B-Instruct-v0.3 on NQ-open dataset. This subset includes 500 queries that can be answered directly using parametric knowledge (Known) and 500 queries that require external knowledge for answering (Unknown). We then retrieve 100 documents for each query from the Wikipedia dump to serve as the grounding data. Overall, we have a total of 100,000 original instances for the following perturbation step.
Automated Perturbation
As introduced in Section 3.1, the perturbation $g(.)$ injects spurious features by modifying the grounding data. For style and source perturbations, $g(.)$ is implemented using an LLM [^2] Unless otherwise specified, all model-based $g(.)$ are implemented using Llama-3.1-70B-Instruct. prompted by carefully crafted guidelines to modify the raw document, producing counterparts of the original instances. For logic and format perturbations, we develop $g(.)$ as a heuristic method based on a set of predefined rules [^3] One exception is that we implement the LLM-reranked perturbation using an LLM-based $g(.)$.. To simulate metadata that may appear in the real world, we first synthesize pseudo Wikipedia or Twitter links for the raw instances, and then organize them into HTML format using a rule-based $g(.)$. The complete implementation details for automated perturbation are provided in Appendix C.
3.4 Causal Features Preservation
To eliminate the effect of causal features, it is essential to follow the principle of controlled experiments by keeping causal features constant while systematically manipulating spurious features. This approach isolates the impact of spurious features from that of causal features, enabling an accurate quantification of robustness against spurious features. In our framework, we introduce two methods to ensure the stability of causal features in the grounding data.
Maintain Semantic Equivalence
For models capable of following human instructions, we directly instruct them to maintain semantic equivalence when injecting spurious features. Nonetheless, it’s impossible to completely avoid semantic shift during the perturbation process. To ensure the semantic consistency before and after introducing perturbation, we employ a bidirectional entailment algorithm to filter out instance pairs (raw instance, perturbed instance) with semantic inequivalence. Specifically, for document $G$ and its modified counterpart $g(G)$, we use a Natural Language Inference (NLI) system to detect whether the latter can be inferred from the former, and vice versa. The NLI system classifies predictions into one of: entailment, neutral, contradiction. We compute both directions, and the algorithm returns equivalent if and only if both directions are predicted as entailment.
In general, this algorithm can be implemented by any NLI system. However, in our case, the concatenation of $G$ and $g(G)$ sometimes exceeds the context limitation of a Bert-based NLI model. Hence, we apply an LLM-based NLI system  [^4] Farquhar et al. (2024) confirms the effectiveness of the LLM-based NLI system through human annotation, demonstrating that its performance is on par with the DeBERTa-large model used in Kuhn et al. (2023). to implement the bidirectional entailment algorithm. The prompt is included in Appendix D.
Preserve Ground Truths
While semantic equivalence protects causal features to the greatest extent, the perturbation may lead to the correct answer being paraphrased into an alias (e.g., ”President Roosevelt” to ”Roosevelt”). These variations in the grounding data are likely to result in false negatives when determining response correctness, despite the NQ-Open dataset providing multiple potential answer variants for each query. To address this issue, we employ a simple string-matching strategy to filter out documents that undergo unexpected modifications. For Golden documents that originally contained the correct answers, we keep them only if they preserve the ground truths after perturbation. For Noise documents that did not contain the correct answers, we discard them if they unexpectedly acquire the ground truths due to perturbations.
3.5 Overview of the Synthetic Dataset
Through the steps of spurious features injection and causal features preservation, we derive the final dataset available for robustness evaluation. The synthetic dataset generated by the SURE framework is divided into four subsets based on the categories of queries and documents within the instances. Notably, the distribution of the dataset is model-specific, as the classification of Known and Unknown queries is determined by the intrinsic knowledge of the target LLM. Table 1 presents the dataset statistics for assessing Mistral-7B-Instruct-v0.3. And the distribution for Llama-3.1-8B-Instruct is shown in Appendix E.
K-G K-N U-G U-N Total
Style 7766 31152 2593 37692 79203
Source 9249 32435 3228 39101 84013
Logic 9724 35537 3587 41990 90838
Format 11037 38018 4141 45518 98714
Meta 11104 38018 4255 45420 98797
Table 1: Statistics of the evaluation dataset for Mistral-7B-Instruct-v0.3. K-G denotes the instances composed of (Known query, Golden Document), while U-N refers to the instances consisting of (Unknown query, Noise Document). The values represents the number of instance pairs for each type of perturbations within the category of spurious features.
3.6 Robustness Evaluation
We employ an evaluation method $Y(.)$, in line with  Liu et al. (2024); Cuconasu et al. (2024), to measure the correctness of responses generated by RAG systems. This approach checks whether any of the correct answers is contained within the response produced by the LLM and then derives a binary label. Previous researches use accuracy as the primary metric and report it at dataset level to assess the robustness of RALMs, which is quantified by calculating the variations in the models’ accuracy across different types of noise. However, dataset-level metrics has certain limitations, as it may fail to capture fine-grained variations that occur at the instance level. As shown in Figure 3, RALMs may appear robust at dataset-level evaluations but exhibit significant sensitivity at the instance level.
To quantify whether a RAG system is robust and unbiased at the instance level, we assign a ternary label to each instance by comparing the correctness of the LLM’s response before and after introducing the perturbation. This comparison process can be formulated as $C=Y(y_{i})-Y(\hat{y_{i}})$, where $C$ lies in the set $(-1,0,1)$. Based on the comparison outcomes, we define three metrics: Robustness Rate (RR), Win Rate (WR), and Lose Rate (LR). The RR is calculated as follows:
$\text{RR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(C==0)$ (5)
where $N$ is the total number of instances in the dataset; $y_{i}$ and $\hat{y_{i}}$ represent the outputs of LLM for the original and perturbed instances. RR measures the proportion of instances where the RALM’s answer remains consistent (0) before and after introducing the perturbation. Similarly, WR and LR quantify the proportions of instances where the correctness of the RALM’s response changes after the perturbation, either from incorrect to correct ($C==-1$) or from correct to incorrect ($C==1$).
Figure 3: A comparison of dataset-level metric (Acc) and instance-level metric (RR) for robustness evaluation. ✔and ✗indicate the correctness of responses. In this example, the value of RR reflects the unrobustness at the instance level, while dataset-level metric overlook the sensitivity of RALMs to spurious features within documents.
4 Experiments
Mistral-7B-Instruct-v0.3
Spurious Features Perturbations Known-Golden Known-Noise Unknown-Golden Unknown-Noise Total
LR RR WR LR RR WR LR RR WR LR RR WR LR RR WR
Style Simple 7.33 85.00 7,67 4.45 91.64 3.90 7.87 82.95 9.18 0.70 98.76 0.54 3.06 94.09 2.85
Complex 6.05 87.42 6.53 3.85 92.03 4.12 6.90 85.92 7.17 0.62 98.82 0.56 2.63 94.61 2.76
[…]
Self-Generated 6.30 87.06 6.64 3.94 92.02 4.04 6.26 86.80 6.94 0.61 98.77 0.62 2.74 94.42 2.85
[…]
Random 4.47 91.87 3.66 2.43 95.15 2.42 4.18 91.44 4.38 0.36 99.27 0.37 1.76 96.56 1.68
LLM-Ranked 3.52 93.15 3.33 2.07 95.84 2.09 3.57 92.89 3.54 0.34 99.30 0.36 1.48 97.04 1.48
[…]
HTML 9.30 87.03 3.67 5.89 92.36 1.74 8.36 87.39 4.25 0.74 99.01 0.26 4.00 94.62 1.38
YAML 4.75 90.90 4.35 3.88 93.24 2.87 5.05 90.53 4.42 0.51 99.06 0.44 2.47 95.55 1.98
Markdown 3.98 92.49 3.53 2.91 94.36 2.72 4.11 92.59 3.31 0.44 99.15 0.41 1.94 96.29 1.77
[…]
Timestamp (post) 2.74 94.87 2.40 1.16 97.63 1.21 3.45 94.41 2.14 0.17 99.68 0.15 0.98 98.12 0.89
Datasource (wiki) 3.78 92.31 3.91 1.5 96.66 1.84 3.69 92.95 3.36 0.26 99.48 0.26 1.28 97.31 1.41
Datasource (twitter) 2.68 93.59 3.73 1.3 97.22 1.48 2.04 94.90 3.06 0.20 99.59 0.21 0.98 97.80 1.22
Table 2: Robustness evaluation results of Mistral-7B-Instruct-v0.3 on the synthetic dataset. For timestamp perturbation, (pre) and (post) represents whether the virtual timestamp we created is before or after the LLM’s knowledge cutoff date. For datasource perturbation, (wiki) and (twitter) indicate the domains of the pseudo link we synthesize. We use Bold to mark the WR values that are higher than the LR, suggesting that the perturbation is beneficial.
4.1 Experimental Setup
We assess the robustness of RAG systems to spurious features by evaluating them on their most popular application—the Question Answering (QA) task, following the standard ”retrieve-read” setting of the RAG paradigm. We employ Contriever-msmarco as the default retriever and conduct our main experiments using the synthetic dataset generated by the SURE framework. Due to space limitations, additional implementation details can be found in Appendix F.
4.2 Result Analysis
In this section, we analyze our main results by category and offer insights from various perspectives. The results of Mistral-7B-Instruct and Llama-3.1-8B-Instruct are presented in Table 2 and Table 6, respectively.
For Different Queries
There is a significant difference in robustness rates between known and unknown queries when evaluated on noise documents (compare K-N and U-N), whereas no such gap is observed for golden documents (compare K-G and U-G). This phenomenon arises from the tug-of-war between an LLM’s internal prior and external evidence. When the retrieved content is noise documents, LLMs are likely to override it with their own correct prior knowledge (Known). However, the robustness to explicit noise may be affected by the spurious features in documents. In other words, implicit noise (spurious features) can coexist with explicit noise, highlighting the challenge of maintaining robustness to spurious features in complex noise scenarios.
For Different Grounding Data
Across all types of spurious features, the robustness rate on noise documents is consistently higher than that on golden documents, whether for known or unknown queries. This derives from the limitation of the RR metric we propose, which measures unrobustness by capturing changes in answer correctness rather than the inconsistency of responses. When tested on noise documents, regardless of which spurious features are injected, LLMs always generate incorrect responses, as noise documents lack ground truths. In this case, even though the responses change, the RR does not decrease since all responses remain incorrect. However, in real-world applications, our primary concern is whether the RAG system can consistently generate correct answer when faced with golden documents containing different spurious features. Therefore, we primarily focus on the RR results for the (K, G) and (U, G) subsets.
For Different Metrics
If the win rate surpasses the lose rate, it shows that more instances were corrected rather than misanswered after introducing perturbations. Based on the comparison of WR and LR in Table 2 and Table 6, we can conclude that Not every spurious feature is harmful and they can even be beneficial sometimes. For example, the WR of source perturbations is consistently higher than the LR on both models. This suggests that the performance of RAG can be enhanced by introducing beneficial spurious features (e.g., simply paraphrasing documents using LLMs).
For Different Perturbations
We observe notable differences in the robustness rates among five types of spurious features. However, within each category, the robustness rates across different perturbations are relatively close. When further comparing perturbations within the same category, we find that while their RR values are comparable, their WR and LR can differ significantly, indicating LLMs’ preference for certain sub-perturbations. In summary, there are variations in robustness across different spurious features, and preference exists among the sub-perturbations of each spurious feature.
4.3 SIG Benchmark & Further Analysis
The raw synthetic dataset is not ideal for extensive evaluation due to its large size. Furthermore, the class imbalance result in unfair comparisons across different types of spurious features. To facilitate more efficient evaluation, we extract the most challenging data from the subsets of synthetic dataset to create a lightweight benchmark: SIG (Spurious features In Golden document). Specifically, we select instance pairs where both the Mistral and Llama models exhibit sensitivity and unrobustness in our main experiments. An equal number of samples (100) are chosen from the (K,G) and (U,G) subsets of each perturbations.
Robustness Comparison of SOTA LLMs
We evaluate a broader range of models on SIG benchmark, including GPT-4O, GPT-4O-mini, Mistral-Large-Instruct  [^5] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama-3.3-70B-Instruct, Deepseek-v3 (671B,MoE), and Qwen2.5-72B-Instruct. To better compare the robustness of different models, we average the RR of each perturbation within a category to derive the overall robustness for a specific type of spurious feature. The performance of six SOTA LLMs is then visualized using a radar chart, as shown in Figure 4. Notably, GPT-4O-mini achieved the best performance, even surpassing GPT-4O by a large margin on the format and meta types. Despite the impressive robustness of closed-source models, they may still exhibit sensitivity to certain specific perturbations. For instance, GPT-4o achieved only an 89% robustness rate on the datasource(twitter) perturbation.
Figure 4: Comparison of Robustness rates across six SOTA LLMs.
Scaling Analysis for Different Model Sizes
To investigate the impact of parameter scale on RAG robustness, we gradually increase the size of LLM-based readers (Qwen2.5 series, ranging from 0.5B to 72B) and evaluate their robustness across five types of spurious features. As illustrated in Figure 5, the robustness rate for all spurious features shows a relatively upward trend as the model size increases. However, when we further scale the model from 32B to 72B, the RR undergoes a significant decline (except for format and meta). This indicates that robustness issues related to spurious features cannot be resolved simply by increasing model size. Interestingly, for meta perturbations, while RALMs demonstrate strong robustness across all scales (even for the 0.5B model), their performance receives little to no benefit from scaling up.
Figure 5: Scaling analysis of robustness to spurious features.
5 Conclusion
In this work, we formally highlight the spurious features problem in RAG system. Through preliminary experiments, we provide statistical evidence to support the presence of spurious features in RALMs. We also propose a novel evaluation framework, SURE, to assess the robustness of LLMs against spurious features. This framework includes a comprehensive taxonomy of spurious features, carefully designed metrics for robustness evaluation, and a data synthesis pipeline. Extensive evaluation and further analysis offers valuable insights for future research.
Impact Statement
We introduce a new type of noise, spurious features, that can undermine the robustness and trustworthiness of contemporary RAG systems. With the increasing diversity of golden documents on the internet, this noise may become more prevalent. Furthermore, traditional methods for improving the robustness of RAG systems are ineffective for handling spurious features, as they are implicit and can coexist with golden documents. Therefore, harmful spurious features could undermines user trust and contaminates Internet data. Moving forward, we aim to explore methods for mitigating hallucinations and robustness issues caused by spurious features within grounding data.
Appendix A Related Work
A.1 Robustness Evaluation of Retrieval-Augmented Generation
RAG systems comprise two core components: a retriever and an LLM-based reader. Augmenting LLMs with retrieved external knowledge has been proven to effectively reduce hallucinations (Shuster et al., 2021; Kang et al., 2023). However, the retrieved contexts inevitably contains noise in addition to desirable knowledge, which may mislead LLMs to produce an incorrect response (Bian et al., 2024; Feldman et al., 2024). Previous works have explored automated evaluation frameworks to assess the robustness of RAG systems in various settings. For instance, Chen et al. (2024a) benchmarked four fundamental capabilities required for RAG, including noise robustness, negative rejection, information integration and counterfactual robustness. Some studies have provided a detailed taxonomy of noise documents to further simulate the complexity of real-world scenarios and highlighted the potential positive effects of certain types of noise (Cuconasu et al., 2024; Wu et al., 2024a). There are also some recent works that propose using LLM-as-a-judge (Li et al., 2024a) to evaluate the RAG system (Wang et al., 2024b).
While these studies have identified several explicit noises that affect the robustness of RAG systems, they overlook implicit noises. This type of noise, such as phrasing and formatting, is everywhere and unavoidable, as it coexists with the grounding data without altering its semantic information. In this work, we define these semantic-agnostic noises as spurious features and evaluate the robustness of RALMs to such noises.
A.2 Prompt Sensitivity of LLMs
Large Language Models take prompts as inputs and then generate response accordingly. Prompts are instructions provided to an LLM to perform specific tasks automatically and ensure desired qualities in the generated output. However, it is known that current LLMs are sensitive to the features of input prompts (Zhu et al., 2023). This sensitivity poses challenges for researchers attempting to evaluate the model’s performance accurately and precisely (Zhuo et al., 2024).
Some existing works have investigated the impact of different prompt techniques on model performance, including chain-of-thought (Wei et al., 2022), in-context learning (Min et al., 2022), and role-play prompting (Kong et al., 2024). Beyond these causal features that significantly influence the meaning of prompts, other works have demonstrated that LLMs are highly sensitive to spurious features (Sclar et al., 2024), e.g, prompt formatting (He et al., 2024), language style (Li et al., 2023a), the order of options (Pezeshkpour & Hruschka, 2024).
Currently, there is no statistical or empirical evidence to support the existence of spurious features in RALMs. To address this gap, we extend the definition of spurious features to RAG systems through statistical testing and empirical analysis.
Appendix B Preliminary Experiment Results
We introduce several semantic-agnostic features we measured in our preliminary experiments. These features include:
  • Flesch Score: A readability metric designed to evaluate text difficulty. It is calculated based on the average number of syllables per word and the average number of words per sentence. The Flesch score is a number on a scale from 0 to 100, where a higher score indicates that the text is easier to read.
  • Distinct-1: A metric used to assess the diversity of generated text. It calculates the proportion of unique words (distinct words) to the total number of words in the output. A higher Distinct-1 score indicates that the text contains a greater variety of unique words, implying more diversity in the generated content.
  • Dependency Tree Depth (DTD): A syntactic complexity metric calculated by analyzing its dependency tree. Dependency Tree Depth refers to the maximum depth of a sentence’s dependency parse tree. A deeper tree suggests more complex sentence structures, while a shallower tree indicates simpler syntactic constructions.
  • Perplexity (PPL): A metric used for evaluating language models, measuring how well a probabilistic model predicts a given text. It reflects the uncertainty of a language model when generating sequences of words. Lower PPL values indicate better predictive performance, meaning the model assigns higher probabilities to the actual labels in the sequence.
  • Token Length: We compute the total number of tokens in a text as an alternative measure of text length, given that the documents in our corpus have been pre-segmented into fixed 100-word chunks. The value is model-specific and depends on the model’s vocabulary.
Kolmogorov-Smirnov (K-S) Test
The K-S test is a non-parametric statistical test used to compare the distribution of two datasets. It evaluate whether two samples come from the same underlying probability distribution. The null hypothesis of the K-S test is that the two samples are drawn from the same distribution, while the alternative hypothesis is that the two samples are drawn from different distributions. There are two key values provided by K-S test: the K-S Statistic quantifies the largest difference between the two sample distributions, and the p-value assess the statistical significance of that difference. If the p-value is lower than a chosen significance level (0.05 in our experiments), we reject the null hypothesis, concluding that the two distributions are significantly different. Otherwise, we fail to reject the null hypothesis, suggesting that there is no significant difference between the two distributions. We provide the K-S test results in Table 3 and Table 4, with the corresponding distribution visualizations plots shown in Figure 6 and Figure 7.
Experimental Group Control Group
K-S statistic P-value K-S statistic P-value
[…]
Table 3: K-S test results for Mistral-7B-Instruct-v0.3 as the oracle retriever.
Experimental Group Control Group
K-S statistic P-value K-S statistic P-value
Flesch score 0.0305 0.1694 0.0173 0.8210
[…]
DTD 0.0474 0.0051∗∗ 0.0203 0.6431
PPL 0.0538 0.0009∗∗∗ 0.0181 0.7791
[…]
Table 4: K-S test results for Llama-3.1-8B-Instruct as the oracle retriever.
(a) Feature distributions of the experimental group
(b) Feature distribution of the control group
Figure 6: Visualization of feature distributions for Mistral-7B-Instruct-v0.3
(a) Feature distributions of the experimental group
(b) Feature distribution of the control group
Figure 7: Visualization of feature distributions for Llama-3.1-8B-Instruct
Appendix C Implementation Details for Injecting Spurious Features
We provide detailed prompts for LLM-based perturbations in Figure 8. For rule-based perturbations, placeholder template is presented in Figure 9.
Style Perturbations [Simple] Please simplify the following text while preserving its original meaning. Use shorter sentences, basic vocabulary, and clear language. Avoid complex structures, technical terms, or ambiguous expressions. Here is the passage to simplify:{Document} [Complex] Please complexify the following text while preserving its original meaning. Use longer sentences, intricate sentence structures, and advanced vocabulary. Avoid contractions, informal language, and colloquial expressions, ensuring the text maintains a professional and authoritative tone throughout. Here is the passage to complexify:{Document} Source Perturbations Please rewrite the following passage. Ensure that the overall meaning, tone, and important details remain intact. Avoid any significant shifts in style or focus. The aim is to create a fresh version while faithfully conveying the original content. Here is the passage to paraphrase:{Document} Logic Perturbations [LLM-Ranked] Rearrange the following list of sentences in your preferred logical order and provide only the indices of the sentences. Please do not include any explanations. Example:{Example} Sentences List:{Sentences List} The length of the Sentences List is {Length of Sentences List}. Therefore, the indices must contain {Length of Sentences List} elements, and the index values cannot exceed {Length of Sentences List - 1}. [Reverse] [Pyhton Code] [Random] [Python Code]
Figure 8: Prompt templates for LLM-based perturbations.
Format Perturbations [JSON]
    "title": "{Title}",
    "text": "{Document}"
[HTML]
<html lang="en">
<head>
    <meta charset="UTF-8">
    {Title}
</head>
<body> {Document} </body>
</html>
[YAML]
Title: {Title}
Text: {Document}
[Markdown]
# {Title}
{Document}
Metadata Perturbations [Timestamp]
<html lang="en">
<head>
    <meta charset="UTF-8">
    <meta name=’timestamp’ content=’{timestamp}’>
    {Title}
</head>
<body> {Document} </body>
</html>
[Datasource]
<html lang="en">
<head>
    <meta charset="UTF-8">
    <meta name=’datasource’ content=’{datasource}’>
    {Title}
</head>
<body> {Document} </body>
</html>
Figure 9: Placeholder templates for rule-Based perturbations.
Appendix D Implementation Details for Preserving Causal Features
We employ a bidirectional entailment algorithm to ensure the semantic equivalence before and after introducing spurious features. The prompts for its core component, NLI model, are shown in Figure 10.
Consider the two passages below. Premise: {raw text} Hypothesis: {perturbated text} Does the premise semantically entail the hypothesis? Answer with ’entailment’ if they are paraphrases,’contradiction’ if they have opposing meanings, or ’neutral’ if they are neither. Response:
Figure 10: Prompts for LLM-based NLI system.
Appendix E Statistics of the Synthetic Dataset
We present the dataset statistics for evaluating Llama-3.1-8B-Instruct in Table 5.
K-G K-N U-G U-N Total
Style 7321 28975 3038 39869 79203
Source 8768 30145 3709 41391 84013
Logic 9229 33294 4082 44233 90838
Format 10481 35616 4697 47920 98714
Meta 10563 35451 4796 47987 98797
Table 5: Distribution of the synthetic dataset for Llama-3.1-8B-Instruct.
Appendix F Experimental Setup Details
Dataset
We conduct our main experiments using the synthetic dataset generated by the SURE framework. The queries are sourced from the NQ-open dataset  [^6] https://huggingface.co/datasets/google-research-datasets/nq_open, and the documents are derive from the English Wikipedia dump.
Models
We test two representative LLMs in our main experiments: Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct.
Prompts
The instruction $I$ in the RAG prompt $P=(I,G,Q)$, shown in Figure 11, is derived from Cuconasu et al. (2024), with slight modifications to better adapt to our setting.
Implementation Details
We follow the typical ”retrieve-read” setting of RAG paradigm. For the retrieval module, we use Contriever-msmarco [^7] https://huggingface.co/facebook/contriever-msmarco, a BERT-based dense retriever, as the default retriever. It is finetuned on the MS MARCO dataset (Bajaj et al., 2016) after unsupervised pretraining via contrastive learning (Izacard et al., 2021). To optimize the efficiency of vector similarity searches, we employ the Faiss library (Douze et al., 2024). For the read module, we deploy LLMs on NVIDIA A100 GPUs and accelerate inference with vllm [^8] https://github.com/vllm-project/vllm. We follow (Li et al., 2023b; Zhang et al., 2024a, b) in using a greedy decoding strategy and set the temperature to 0.1 to ensure stable outputs and strong reproducibility.
You are given a question and you MUST respond by EXTRACTING the answer (max 5 tokens) from the provided document. If the document does not contain the answer, respond with NO-RES.
Figure 11: Instruction $I$ used for the QA task.
Llama-3.1-8B-Instruct
Spurious Features Perturbations Known-Golden Known-Noise Unknown-Golden Unknown-Noise Total
LR RR WR LR RR WR LR RR WR LR RR WR LR RR WR
Style Simple 7.79 83.04 9.18 1.70 95.80 2.50 8.43 82.88 8.69 0.27 99.45 0.28 1.80 95.96 2.24
Complex 6.00 85.60 8.40 1.91 96.59 1.50 6.71 84.86 8.43 0.23 99.57 0.20 1.63 96.62 1.75
Source LLM-Generated 5.89 86.43 7.69 1.43 96.83 1.74 6.20 85.71 8.09 0.22 99.56 0.22 1.51 96.60 1.89
Self-Generated 6.55 85.01 8.44 1.55 96.37 2.09 6.52 86.36 7.12 0.19 99.57 0.24 1.62 96.32 2.06
Logic Reverse 5.06 90.82 4.12 1.13 97.82 1.06 5.73 89.71 4.56 0.21 99.67 0.13 1.29 97.64 1.07
Random 3.91 93.16 2.93 0.86 98.31 0.83 4.21 91.67 4.12 0.14 99.72 0.14 0.97 98.18 0.85
LLM-Ranked 3.24 93.93 2.83 0.82 98.43 0.74 3.58 93.36 3.06 0.13 99.76 0.11 0.85 98.39 0.75
Format JSON 7.01 88.25 4.74 1.70 97.25 1.05 5.92 89.63 4.45 0.25 99.61 0.14 1.76 97.08 1.16
HTML 11.85 84.46 3.69 2.70 96.90 0.40 9.33 86.78 3.90 0.35 99.61 0.04 2.85 96.41 0.74
YAML 5.26 89.94 4.80 1.26 97.41 1.33 4.79 90.80 4.41 0.17 99.67 0.16 1.32 97.40 1.28
Markdown 2.32 92.23 5.45 0.60 96.89 2.51 2.34 93.46 4.19 0.07 99.61 0.32 0.61 97.55 1.84
Metadata Timestamp (pre) 2.08 95.81 2.11 0.28 99.42 0.29 2.54 95.56 1.90 0.02 99.95 0.03 0.46 99.10 0.44
Timestamp (post) 2.04 95.86 2.10 0.25 99.43 0.32 2.81 95.56 1.63 0.02 99.95 0.03 0.46 99.11 0.43
Datasource (wiki) 2.11 93.45 4.44 0.23 98.96 0.81 3.25 92.47 4.27 0.03 99.86 0.11 0.48 98.50 1.03
Datasource (twitter) 2.27 94.11 3.62 0.31 99.25 0.43 2.77 93.97 3.25 0.02 99.91 0.07 0.50 98.77 0.73
Table 6: Robustness evaluation results of Llama-3.1-8B-Instruct on the synthetic dataset. We use Bold to mark the WR values that are higher than the LR, suggesting that the perturbation is beneficial.

### v2

@@
-Quantifying the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data
+Quantifying and Improving the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data
 Abstract
 Robustness has become a critical attribute for the deployment of RAG systems in real-world applications.
-Existing research focuses on robustness to explicit noise (e.g., document semantics) but overlooks spurious features (a.k.a.
-implicit noise).
-While previous works have explored spurious features in LLMs, they are limited to specific features (e.g., formats) and narrow scenarios (e.g., ICL).
-In this work, we statistically confirm the presence of spurious features in the RAG paradigm, a robustness problem caused by the sensitivity of LLMs to semantic-agnostic features.
-Moreover, we provide a comprehensive taxonomy of spurious features and empirically quantify their impact through controlled experiments.
-Further analysis reveals that not all spurious features are harmful and they can even be beneficial sometimes.
-Extensive evaluation results across multiple LLMs suggest that spurious features are a widespread and challenging problem in the field of RAG.
-We release all codes and data at: https://github.com/maybenotime/RAG-SpuriousFeatures.
-Machine Learning, ICML
+Existing research focuses on robustness to explicit noise (e.g., document semantics) but overlooks implicit noise (spurious features).
+Moreover, previous studies on spurious features in LLMs are limited to specific types (e.g., formats) and narrow scenarios (e.g., ICL).
+In this work, we identify and study spurious features in the RAG paradigm, a robustness issue caused by the sensitivity of LLMs to semantic-agnostic features.
+We then propose a novel framework, SURE, to empirically quantify the robustness of RALMs against spurious features.
+Beyond providing a comprehensive taxonomy and metrics for evaluation, the framework’s data synthesis pipeline facilitates training-based strategies to improve robustness.
+Further analysis suggests that spurious features are a widespread and challenging problem in the field of RAG.
+Our code is available at https://github.com/maybenotime/RAG-SpuriousFeatures.
+Quantifying and Improving the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data
+Shiping Yang1 2thanks: Work done during internship, Jie Wu2thanks: Correspondence to: jiewu1@microsoft.com, Wenbiao Ding2, Ning Wu2, Shining Liang2, Ming Gong2, Hongzhi Li2, Hengyuan Zhang3, Angel X.
+Chang1, Dongmei Zhang2 1Simon Fraser University 2Microsoft 3The University of Hong Kong
 1 Introduction
-Figure 1: An example from the SURE framework (Sec.
-3), illustrating the sensitivity of RAG systems to spurious features within grounding data.
-MD represents the Markdown format.
+Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm to mitigate LLMs hallucinations Gao et al.
+(2023); Yang et al.
+(2023), integrating relevant external knowledge to improve the factuality and trustworthiness of LLM-generated outputs Zhou et al.
+(2024).
+However, Retrieval-Augmented Language Models (RALMs) [^1] We focuses on the generation end of RAG systems (i.e., RALMs), and hereafter RAG systems refer to RALMs.
+still face substantial robustness issue due to the presence of noise in retrieved documents Liu et al.
+(2023); Li et al.
+(2024).
+Figure 1: An example from the SURE_Wiki dataset (Sec.
+4), illustrating the sensitivity of RAG systems to spurious features within grounding data.
 The original retrieved document is fed into the LLMs in different formats, leading to inconsistent responses.
-Retrieval-Augmented Generation (RAG) has emerged as a promising paradigm to mitigate LLMs hallucinations (Gao et al., 2023; Yang et al., 2023b), integrating relevant external knowledge to improve the factuality and trustworthiness of LLM-generated outputs (Zhou et al., 2024).
-However, Retrieval-Augmented Language Models (RALMs) still face substantial robustness issue due to the presence of noise in retrieved documents (Liu et al., 2023; Li et al., 2024c).
-Recent research aims to explore the characteristics that affect the robustness of RAG systems from the perspective of grounding data construction (Cuconasu et al., 2024).
-These studies examine various factors, including the type (Wu et al., 2024a), number (Xu et al., 2024), and position of documents (Liu et al., 2024) within the prompt context.
-A more detailed discussion of related work is in Appendix A.
-However, previous analyses primarily focus on explicit noise (i.e., causal features) that significantly alter the semantic information of grounding data(Wu et al., 2024b; Cuconasu et al., 2024), while neglecting spurious features (a.k.a.
-implicit noise) that make slight modifications but still preserve the meaning of contexts.
-This limitation extends to existing evaluation benchmarks, which simulate complex noise scenarios to assess the robustness of RAG systems (Chen et al., 2024a; Wu et al., 2024a), yet lack available benchmarks and metrics to measure the robustness of LLMs against spurious features.
-Contemporary RAG systems typically employ production-level retrievers, such as Bing and Google, to collect relevant information from the internet.
+Recent research aims to explore the characteristics of grounding data that influence the robustness of RAG systems Cuconasu et al.
+(2024).
+These studies examine various factors, including the type Wu et al.
+(2024a), number Xu et al.
+(2024), and position of documents Liu et al.
+(2024) within the prompt context.
+However, previous analyses primarily focus on explicit noise that significantly alter the semantic information (causal features) of grounding data Wu et al.
+(2024b); Cuconasu et al.
+(2024), while neglecting implicit noise (spurious features) that introduce semantic-agnostic modifications.
+This limitation extends to existing evaluation benchmarks, which simulate complex noise scenarios to assess the robustness of RAG systems Chen et al.
+(2024a); Wu et al.
+(2024a), yet lack available benchmarks and metrics to measure the robustness of RALMs against spurious features.
+Contemporary RAG systems typically employ production-level retrievers, such as Bing and Google, to retrieve relevant information from the internet.
 Unlike a single corpus, the internet encompasses diverse data with distinct features.
@@
 (2024), which demonstrate that LLMs are extremely sensitive to the format of prompts (i.e., spurious features).
-For more related work, see Appendix A.
-Unfortunately, there is no statistic and empirical evidence to support the existence of spurious features in the RAG paradigm.
-This highlights the urgent need to redefine spurious features in RAG and systematically quantify the robustness of RALMs against them.
-To address these challenges, we first design a preliminary experiment to demonstrate that RALMs are sensitive to semantic-agnostic features in the grounding data, thereby extending the definition of spurious features to RAG systems.
-Building on findings from our preliminary experiment and recent studies, we identify five types of spurious features that may appear in RAG scenarios.
-Then, we propose a novel framework, SURE, for automating the process of robustness evaluation.
+However, spurious features in the RAG paradigm have not been empirically and systematically studied, nor have effective mitigation strategies been proposed.
+In this work, we first propose a novel framework, SURE, for automating the process of robustness evaluation.
 This framework follows a perturb-then-evaluate approach, offering great scalability.
-In SURE, automated perturbations are applied on the original instances to inject the corresponding spurious features.
+In SURE, automated perturbations are applied to the original instances to inject the corresponding spurious features.
 The perturbed instances are then examined to ensure that the causal features remain intact.
 After these steps, we employ tailored metrics to quantify the robustness of RALMs against spurious features.
-Further analysis reveals that not every spurious features is harmful and they can even be beneficial sometimes.
-We also find that our proposed metrics are more suitable for evaluating spurious features in golden documents.
-Based on this finding, we distill the most challenging instances from the synthetic data generated by our framework to create a lighter benchmark, SIG, enabling more efficient robustness evaluation.
-We evaluate 12 representative LLMs with varying architectures and scales, and the results show that maintaining robustness against spurious features remains a significant challenge, especially for open-source LLMs.
-Our contribution can be summarized as follows: 1) We extend the definition of spurious features to RAG systems through a preliminary experiment.
-To the best of our knowledge, this is the first comprehensive study to define and evaluate spurious features from RAG perspective.
-2) We propose a novel evaluation framework, SURE, to assess the robustness of RALMs against spurious features, which effectively simulate five types of spurious features that may appear in real world scenarios.
-3) Through extensive experiments and analysis using SURE on two representative LLMs, we provide valuable insights for future research.
-4) Based on these findings, we curate a lightweight yet challenging evaluation dataset named SIG and benchmark the robustness of the current state-of-the-art LLMs against spurious features.
-2 Preliminary
-In this section, we first define causal and spurious features in the context of retrieval-augmented generation and then provide statistical evidence to support the existence of spurious features.
-Specifically, we start by defining an oracle retriever to sample data according to the preferences of LLMs and then apply statistical testing to validate that RALMs exhibit biases toward semantic-agnostic features within the grounding data.
-2.1 Causal and Spurious Features in RAG
-In general, causal features are input features that have a direct causal effect on the output of predictive model (Yu et al., 2020).
-Their relationship is rooted in causality, rather than mere statistical correlation.
-When it comes to Large Language Models, the meaning and intent of prompts serve as causal features that directly influence the models’ responses.
-In the context of RAG, causal features refer to the semantic information of grounding data, including whether the correct answer exists, the amount of noise documents in the grounding data, and the content of those noise documents.
-In contrast, spurious features are input features that co-occur with causal features and are erroneously captured by the model (Neuhaus et al., 2023).
-These features exhibit a statistical correlation with the model’s output but lack a causal relationship.
-Recent research has shown that LLMs are sensitive to seemingly trivial features like prompt formatting, thereby extending the definition of spurious features to LLMs (Sclar et al., 2024).
-Similarly, we hypothesize that the semantic-agnostic features of the grounding data can be defined as spurious features in RAG systems.
-However, the conclusion drawn from LLMs may not be applicable to RALMs.
-Unlike LLMs, the input prompts for RALMs incorporate additional dynamic content—grounding data—to augment their output.
-This grounding data is derived from retrieved documents rather than static instructions provided directly by users.
-Therefore, we design a preliminary experiment to validate whether RALMs are sensitive to semantic-agnostic features within the grounding data.
-2.2 Oracle Retriever
-We aim to confirm the existence of spurious features in RAG scenarios, i.e., to demonstrate the sensitivity of RALMs to semantic-agnostic features in grounding data.
-There are some challenges in revealing the sensitivity of RALMs.
-Specifically, When retrieving from a single corpus, it is difficult to mine semantically equivalent counterparts with obvious differences in semantic-agnostic features.
-This is because documents from the corpus often share similar styles and formats.
-With only slight differences in these features, we are difficult to observe significant performance variations using traditional evaluation metrics like accuracy.
-Thus, more fine-grained metrics are required to capture the subtle performance changes.
-Inspired by the use of LLMs as supervision signals for document utility (Izacard et al., 2023; Gan et al., 2024), we propose the oracle score, which measures fine-grained performance through calculating the log probability of generating correct answers based on the given documents.
-The oracle score is defined as follows:
-[…]
-where $x$ is the input prompt for RALMs, including the instruction $I$, grounding data $G$, and query $Q$; $y$ represents the ground truth answer; $\theta$ denotes the model parameters; and $T$ is the total length of the answer sequence.
-For cases with multiple answers, we compute the final score by averaging the corresponding oracle scores across all answers.
-Notably, the oracle score is less effective at estimating the helpfulness of documents when handling queries with long-sequence answers  [^1] In our experiments, we only use queries where the answer sequences is fewer than 5 tokens, ensuring the effectiveness of oracle scores..
-This limitation arises from the growing expansion of the solution space as the answer length increases.
-We further define an oracle retriever that ranks documents according to their oracle scores.
-Since the oracle score is computed using LLMs as supervision signals, the oracle retriever can be regarded as a sampler that reflects the LLMs’ preferences for grounding data.
-With this sampler, we successfully transform the target of demonstrating that RALMs are sensitive to semantic-agnostic features within grounding data into showing that the oracle retriever is biased toward spurious features in documents.
-2.3 Preliminary Experiment & Analysis
-Using the oracle retriever, we recall 100 documents from the Wikipedia dump for each query in the NQ-open dataset.
-However, the computational cost of calculating oracle scores across the entire corpus to select top-ranked documents is prohibitive.
-Therefore, we introduce Contriever-msmarco, a traditional dense retriever, for first-stage retrieval, followed by reranking in descending order based on the oracle scores.
-In addition to reducing the computational load, this initial retrieval also ensures the semantic similarity of retrieved documents.
-To further eliminate the effect of causal features, documents without golden answers are filtered out, ensuring that the remaining documents have roughly consistent causal features.
-We then select the first-ranked and last-ranked documents for each query from the remaining documents, resulting in two sets, each containing 2658 samples.
-By comparing the differences in feature distributions between these two sets, we can assess whether the oracle retriever exhibits bias toward semantic-agnostic features.
-If these two sets do not belong to the same feature distribution, this can be attributed to the oracle retriever’s bias towards semantic-agnostic features during sampling.
-To confirm that this bias is not introduced by the dense retriever in first-stage retrieval, we establish a control group by randomly sampling two documents instead of selecting the first- and last-ranked documents.
-To evaluate whether the two distributions are same, we employ the Kolmogorov-Smirnov (K-S) test.
-The following semantic-agnostic features are measured in the experiments: 1) Flesh Score, 2) Distinct-1, 3) Dependency Tree Depth, 4) PPL, and 5) Token Length.
-A detailed introduction of the K-S test and these features can be found in the Appendix B.
-We conduct experiments using Mistral-7B-Instruct-v0.3 to implement the oracle retriever.
-The K-S statistic and P-value are presented in Table 3.
-Furthermore, we visualize the feature distributions for both the experimental and control groups in Figure 6.
-For all tested features in the experimental group, the K-S test rejects the null hypothesis, concluding that the distribution of the two sets are significantly different.
-In contrast, for the control group, the K-S test fails to reject the null hypothesis.
-The results for Llama-3.1-8B-Instruct are also provided in Appendix B.
-According to these results, we can conclude that RALMs exhibit bias toward spurious features in documents.
-The preliminary experiment provides statistical evidence supporting the existence of spurious features in RAG systems.
-Nevertheless, it does not offer empirical evidence or quantitative analysis.
-Inspired by previous data synthesis studies (Tan et al., 2024b; Tong et al., 2024; Li et al., 2024b; Wang et al., 2024a), we use a data synthesis approach to better control feature variables and quantify the robustness of RALMs against spurious features.
+To enable more efficient evaluation, we select the most challenging instances to create a lighter benchmark, SIG_Wiki.
+Extensive evaluations across diverse LLMs and methods indicate that maintaining robustness against spurious features is a widespread and challenging issue.
+To address this, we introduce two training-based strategies to mitigate the lack of robustness caused by spurious features, leveraging the synthetic data generated by our SURE framework.
+Our contribution can be summarized as follows: 1) To the best of our knowledge, this is the first comprehensive study to evaluate spurious features from RAG perspective.
+We propose a novel evaluation framework, SURE, to assess the robustness of RALMs against spurious features, which includes a comprehensive taxonomy, tailored metrics, and a data synthesis pipeline.
+2) Using the synthetic dataset generated by the SURE framework, we curate a lightweight yet challenging evaluation benchmark, SIG_Wiki, and introduce two effective training-based strategies to improve the robustness of RALMs.
+3) Further analysis offer valuable insights for future research.
+For example, we found that not every spurious features is harmful and they can even be beneficial sometimes.
+2 Related Work
+2.1 Robustness Evaluation of RAG Systems
+The retrieved contexts inevitably contains noise in addition to desirable knowledge, which may mislead LLMs to produce an incorrect response Bian et al.
+(2024); Feldman et al.
+(2024).
+Previous works have explored automated evaluation frameworks to assess the robustness of RAG systems in various settings Chen et al.
+(2024a); Cuconasu et al.
+(2024).
+Wu et al.
+(2024a) provided a detailed taxonomy of noise documents to further simulate the complexity of real-world scenarios and highlighted the potential positive effects of certain types of noise.
+While these studies have identified several explicit noises that affect the robustness of RAG systems, they predominantly overlook implicit noises.
+Even some works evaluate the influence of implicit noise, they are often limited to specific types, such as typos Cho et al.
+(2024) and formatting Tan et al.
+(2024a).
+In this work, we comprehensively study semantic-agnostic noises (i.e, spurious features) in RAG systems.
+2.2 Prompt Sensitivity of LLMs
+Prompts are instructions provided to an LLM to perform specific tasks automatically and ensure desired qualities in the generated output.
+However, it is known that current LLMs are sensitive to the features of input prompts Zhu et al.
+(2023).
+This sensitivity poses challenges for researchers attempting to evaluate the model’s performance accurately and precisely Zhuo et al.
+(2024).
+Beyond causal features that significantly influence the meaning of prompts Wei et al.
+(2022); Kong et al.
+(2024), existing works have demonstrated that LLMs are highly sensitive to spurious features Sclar et al.
+(2024) in non-RAG scenarios, e.g, prompt formatting He et al.
+(2024), language style Li et al.
+(2023), the order of options Pezeshkpour and Hruschka (2024).
+In contrast, our work focuses on spurious features within the context of RAG.
+Unlike non-RAG scenarios (e.g., ICL), where spurious features appear in user-provided instructions and remain static across queries, in RAG they occur in retrieved documents and vary dynamically for each query.
 3 Proposed Framework
@@
 As illustrated in Figure 2, this framework comprise four components: 1) Comprehensive Taxonomy.
-We identify five distinct types of spurious features that may arise in the context of Retrieval Augmented Generation.
+We identify and define five common types of spurious features in RAG scenarios.
 2) Spurious Features Injection.
@@
 4) Robustness Evaluation.
-We report the win, lose and robustness rate at instance level to facilitate a fine-grained evaluation.
+We introduce three metrics (Win Rate, Lose Rate, and Robustness Rate) to facilitate fine-grained, instance-level evaluation.
 3.1 Problem Formulation
@@
 In this work, we use a BERT-based dense retriever to obtain the embedding of query and documents, respectively.
-The relevance score is calculated by computing their dot-product similarity:
-$s(q,d_{i})=E(q)\,\raisebox{-0.86108pt}{$\cdot$}\,E(d_{i}).$ (2)
+The relevance score is calculated by computing their dot-product similarity.
 Then, the Top-k documents with the highest similarity scores are retrieved:
@@
 The outputs of LLM-based reader for $P$ and $\hat{P}$ are compared to evaluate the impact of the introduced perturbation:
-$y=\text{LLM}(P),\quad\hat{y}=\text{LLM}(\hat{P}).$ (4)
+[…]
 3.2 Taxonomy of Spurious Features
-Based on the biased spurious features identified in our preliminary experiment and prior works, we introduce five types of spurious features and their corresponding perturbations in detail as follows.
+We develop a comprehensive taxonomy of spurious features, informed by our preliminary experiments (see Appendix A) and insights from prior research.
+The five types of spurious features and their corresponding perturbations are detailed below.
 Style Perturbations
 The same content can be expressed in different styles, using varying tones, words and sentence structures.
-As shown in Section 2.3, LLMs exhibit biases towards readability-related features.
-Similarly, for humans, the readability of a text can significantly influence its accessibility to the audience (Yang et al., 2023a).
+As shown in Appendix A.2, LLMs exhibit biases towards readability-related features.
 Therefore, we define two perturbations from the perspective of readability style: Simple and Complex.
@@
 Source Perturbations
-LLM-generated content, including both misinformation and correct claims, infiltrates every corner of the internet.
-Recent studies have shown that neural retrievers are biased towards LLM-generated content, leading to the marginalization of human-authored content (Dai et al., 2024; Chen et al., 2024b).
+Recent studies have shown that neural retrievers are biased towards LLM-generated content, leading to the marginalization of human-authored content Dai et al.
+(2024); Chen et al.
+(2024b).
 Moreover, our preliminary experiments demonstrate that LLMs are biased towards the Perplexity (PPL) of text.
@@
 Logic Perturbations
-The arrangement of sentences within a passage typically follows a logical order, ensuring the clarity and coherence in the narrative flow.
-Here, we simulate scenarios where the intrinsic logical chain is disrupted by three different perturbations: Random, Reverse, and LLM-reranked, each representing a distinct sentence ordering strategy.
+In RAG systems, documents are often segmented into multiple chunks and may be retrieved in varying orders.
+Here, we simulate scenarios where the intrinsic logical flaw is disrupted by three different perturbations: Random, Reverse, and LLM-reranked, each representing a distinct sentence ordering strategy.
 Format Perturbations
@@
 These formats are usually processed into plain text before being fed to LLMs.
-To mitigate the loss of structural information during this process, some RAG studies propose using the original format, rather than plain text, to augment the generation (Tan et al., 2024a).
-However, as highlighted in previous research, the prompt format is recognized as a spurious feature that can significantly impact model performance (Sclar et al., 2024; He et al., 2024).
-Therefore, we perturb the original document with four common formats to explore the impact of grounding data format in the context of RAG.
+To mitigate the loss of structural information during this process, some RAG studies propose using the original format, rather than plain text, to augment the generation Tan et al.
+(2024a).
+However, as highlighted in previous research, the prompt format is recognized as a spurious feature that can significantly impact model performance Sclar et al.
+(2024); He et al.
+(2024).
+Therefore, we perturb the original document with four common formats to explore the impact of grounding data format.
 Metadata Perturbations
 Metadata is often included in the HTML results returned by search engines.
-In our framework, we consider two main types of metadata: Timestamp and Data source.
-The timestamp indicates the time when the data was created, and data source identifies the origin of the data.
+In our framework, we focus on two types: Timestamp and Data source.
+The timestamp marks when the data was created, and the data source indicates its origin  [^2] Timestamp may serve as causal features in time-sensitive tasks.
+In our experiments, the NQ dataset generally does not contain time-sensitive queries..
+For timestamp perturbations, pre and post denote whether the timestamp is before or after the LLM’s knowledge cutoff date.
+For data source perturbations, wiki and twitter represent the domains of the URLs.
 3.3 Spurious Features Injection
@@
 An instance is the dynamic component of the prompt $P$, consisting of a query $Q$ and grounding data $G$.
-The queries are drawn from the NQ-open dataset, while our data source is English Wikipedia dump as of 20 December 2018.
-To construct the original instances, we first select 1,000 queries based on the close-book QA results of Mistral-7B-Instruct-v0.3 on NQ-open dataset.
-This subset includes 500 queries that can be answered directly using parametric knowledge (Known) and 500 queries that require external knowledge for answering (Unknown).
-We then retrieve 100 documents for each query from the Wikipedia dump to serve as the grounding data.
-Overall, we have a total of 100,000 original instances for the following perturbation step.
+To construct the original instances, we first select 1,000 queries from the NQ-open dataset.
+For each query, we then retrieve 100 documents from the Wikipedia dump to serve as grounding data, yielding 100,000 instances for the following perturbation step.
 Automated Perturbation
@@
 prompted by carefully crafted guidelines to modify the raw document, producing counterparts of the original instances.
-For logic and format perturbations, we develop $g(.)$ as a heuristic method based on a set of predefined rules [^3] One exception is that we implement the LLM-reranked perturbation using an LLM-based $g(.)$..
-To simulate metadata that may appear in the real world, we first synthesize pseudo Wikipedia or Twitter links for the raw instances, and then organize them into HTML format using a rule-based $g(.)$.
-The complete implementation details for automated perturbation are provided in Appendix C.
+For logic and format perturbations, we develop $g(.)$ as a heuristic method based on a set of predefined rules.
+To simulate real-world metadata, we first synthesize pseudo Wikipedia or Twitter links for the raw instances, and then organize them into HTML format using a rule-based $g(.)$.
+Further details for automated perturbation are provided in Appendix B.
 3.4 Causal Features Preservation
@@
 In our framework, we introduce two methods to ensure the stability of causal features in the grounding data.
+Implementation details can be found in Appendix C.
 Maintain Semantic Equivalence
@@
 to implement the bidirectional entailment algorithm.
-The prompt is included in Appendix D.
 Preserve Ground Truths
-While semantic equivalence protects causal features to the greatest extent, the perturbation may lead to the correct answer being paraphrased into an alias (e.g., ”President Roosevelt” to ”Roosevelt”).
-These variations in the grounding data are likely to result in false negatives when determining response correctness, despite the NQ-Open dataset providing multiple potential answer variants for each query.
-To address this issue, we employ a simple string-matching strategy to filter out documents that undergo unexpected modifications.
-For Golden documents that originally contained the correct answers, we keep them only if they preserve the ground truths after perturbation.
-For Noise documents that did not contain the correct answers, we discard them if they unexpectedly acquire the ground truths due to perturbations.
-3.5 Overview of the Synthetic Dataset
-Through the steps of spurious features injection and causal features preservation, we derive the final dataset available for robustness evaluation.
-The synthetic dataset generated by the SURE framework is divided into four subsets based on the categories of queries and documents within the instances.
-Notably, the distribution of the dataset is model-specific, as the classification of Known and Unknown queries is determined by the intrinsic knowledge of the target LLM.
-Table 1 presents the dataset statistics for assessing Mistral-7B-Instruct-v0.3.
-And the distribution for Llama-3.1-8B-Instruct is shown in Appendix E.
-K-G K-N U-G U-N Total
-Style 7766 31152 2593 37692 79203
-Source 9249 32435 3228 39101 84013
-Logic 9724 35537 3587 41990 90838
-Format 11037 38018 4141 45518 98714
-Meta 11104 38018 4255 45420 98797
-Table 1: Statistics of the evaluation dataset for Mistral-7B-Instruct-v0.3.
-K-G denotes the instances composed of (Known query, Golden Document), while U-N refers to the instances consisting of (Unknown query, Noise Document).
-The values represents the number of instance pairs for each type of perturbations within the category of spurious features.
+While semantic equivalence protects causal features to the greatest extent, the perturbation may lead to the correct answer being paraphrased into an alias (e.g., "President Roosevelt" to "Roosevelt").
+To address this issue, we employ a simple string-matching strategy to filter out documents that have undergone unexpected modifications.
 3.5 Robustness Evaluation
@@
 As shown in Figure 3, RALMs may appear robust at dataset-level evaluations but exhibit significant sensitivity at the instance level.
-To quantify whether a RAG system is robust and unbiased at the instance level, we assign a ternary label to each instance by comparing the correctness of the LLM’s response before and after introducing the perturbation.
-This comparison process can be formulated as $C=Y(y_{i})-Y(\hat{y_{i}})$, where $C$ lies in the set $(-1,0,1)$.
-Based on the comparison outcomes, we define three metrics: Robustness Rate (RR), Win Rate (WR), and Lose Rate (LR).
-The RR is calculated as follows:
+[…]
 $\text{RR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(C==0)$ (3)
@@
 RR measures the proportion of instances where the RALM’s answer remains consistent (0) before and after introducing the perturbation.
-Similarly, WR and LR quantify the proportions of instances where the correctness of the RALM’s response changes after the perturbation, either from incorrect to correct ($C==-1$) or from correct to incorrect ($C==1$).
-Figure 3: A comparison of dataset-level metric (Acc) and instance-level metric (RR) for robustness evaluation.
+Similarly, WR and LR quantify the proportions of instances where the correctness of the RALM’s response changes after the perturbation, either from incorrect to correct (C==1C==-1) or from correct to incorrect ($C==1$).
+Figure 3: A comparison of dataset-level metric (Acc) and instance-level metric ($R$) for robustness evaluation.
 ✔and ✗indicate the correctness of responses.
-In this example, the value of RR reflects the unrobustness at the instance level, while dataset-level metric overlook the sensitivity of RALMs to spurious features within documents.
+In this example, $R$ captures instance-level unrobustness, while Acc overlooks RALMs’ sensitivity to spurious features within documents.
 4 Experiments
 Mistral-7B-Instruct-v0.3
-Spurious Features Perturbations Known-Golden Known-Noise Unknown-Golden Unknown-Noise Total
-LR RR WR LR RR WR LR RR WR LR RR WR LR RR WR
-Style Simple 7.33 85.00 7,67 4.45 91.64 3.90 7.87 82.95 9.18 0.70 98.76 0.54 3.06 94.09 2.85
-Complex 6.05 87.42 6.53 3.85 92.03 4.12 6.90 85.92 7.17 0.62 98.82 0.56 2.63 94.61 2.76
-[…]
-Self-Generated 6.30 87.06 6.64 3.94 92.02 4.04 6.26 86.80 6.94 0.61 98.77 0.62 2.74 94.42 2.85
-[…]
-Random 4.47 91.87 3.66 2.43 95.15 2.42 4.18 91.44 4.38 0.36 99.27 0.37 1.76 96.56 1.68
-LLM-Ranked 3.52 93.15 3.33 2.07 95.84 2.09 3.57 92.89 3.54 0.34 99.30 0.36 1.48 97.04 1.48
-[…]
-HTML 9.30 87.03 3.67 5.89 92.36 1.74 8.36 87.39 4.25 0.74 99.01 0.26 4.00 94.62 1.38
-YAML 4.75 90.90 4.35 3.88 93.24 2.87 5.05 90.53 4.42 0.51 99.06 0.44 2.47 95.55 1.98
-Markdown 3.98 92.49 3.53 2.91 94.36 2.72 4.11 92.59 3.31 0.44 99.15 0.41 1.94 96.29 1.77
-[…]
-Timestamp (post) 2.74 94.87 2.40 1.16 97.63 1.21 3.45 94.41 2.14 0.17 99.68 0.15 0.98 98.12 0.89
-Datasource (wiki) 3.78 92.31 3.91 1.5 96.66 1.84 3.69 92.95 3.36 0.26 99.48 0.26 1.28 97.31 1.41
-Datasource (twitter) 2.68 93.59 3.73 1.3 97.22 1.48 2.04 94.90 3.06 0.20 99.59 0.21 0.98 97.80 1.22
-Table 2: Robustness evaluation results of Mistral-7B-Instruct-v0.3 on the synthetic dataset.
-For timestamp perturbation, (pre) and (post) represents whether the virtual timestamp we created is before or after the LLM’s knowledge cutoff date.
-For datasource perturbation, (wiki) and (twitter) indicate the domains of the pseudo link we synthesize.
+Taxonomy Perturbations Known-Golden Known-Noise Unknown-Golden U-N
+LR $R$ WR Org Acc LR $R$ WR Org Acc LR $R$ WR Org Acc $R$
+Style Simple 7.33 85.00 7.67 73.02 73.37 4.45 91.64 3.90 10.82 10.28 7.87 82.95 9.18 56.31 57.62 98.76
+Complex 6.05 87.42 6.53 73.50 3.85 92.03 4.12 11.10 6.90 85.92 7.17 56.58 98.82
+[1pt/1pt] Source LLM-Generated 5.91 87.62 6.47 71.81 72.36 3.57 92.27 4.16 10.79 11.38 6.41 86.52 7.06 54.46 55.11 98.75
+Self-Generated 6.30 87.06 6.64 72.15 3.94 92.02 4.04 10.89 6.26 86.80 6.94 55.14 98.77
+[1pt/1pt] Logic Reverse 5.44 89.34 5.22 69.91 69.69 2.99 94.10 2.92 11.77 11.70 5.97 88.54 5.49 50.26 49.79 99.04
+Random 4.47 91.87 3.66 69.10 2.43 95.15 2.42 11.76 4.18 91.44 4.38 50.46 99.27
+LLM-Ranked 3.52 93.15 3.33 69.72 2.07 95.84 2.09 11.79 3.57 92.89 3.54 50.24 99.30
+[1pt/1pt] Format JSON 7.96 88.53 3.51 70.81 66.35 5.15 92.68 2.17 10.98 8.00 6.95 88.92 4.13 53.32 50.50 99.02
+HTML 9.30 87.03 3.67 65.18 5.89 92.36 1.74 6.83 8.36 87.39 4.25 49.22 99.01
+YAML 4.75 90.90 4.35 70.41 3.88 93.24 2.87 9.97 5.05 90.53 4.42 52.69 99.06
+Markdown 3.98 92.49 3.53 70.36 2.91 94.36 2.72 10.79 4.11 92.59 3.31 52.52 99.15
+[1pt/1pt] Metadata Timestamp (pre) 2.62 94.90 2.48 65.04 64.90 1.28 97.61 1.11 6.83 6.66 3.15 94.45 2.40 48.08 47.33 99.67
+Timestamp (post) 2.74 94.87 2.40 64.70 1.16 97.63 1.21 6.88 3.45 94.41 2.14 46.77 99.68
+Datasource (wiki) 3.78 92.31 3.91 65.17 1.5 96.66 1.84 7.16 3.69 92.95 3.36 47.76 99.48
+Datasource (twitter) 2.68 93.59 3.73 66.08 1.3 97.22 1.48 7.00 2.04 94.90 3.06 49.10 99.59
+Table 1: Robustness evaluation results of Mistral-7B-Instruct-v0.3 on the SURE_Wiki dataset.
+Org indicates the accuracy on original instances, while Acc refers to the accuracy after introducing perturbations.
 We use Bold to mark the WR values that are higher than the LR, suggesting that the perturbation is beneficial.
+In this section, we assess the robustness of RAG systems to spurious features by evaluating them on their most popular application—the Question Answering (QA) task, following the standard "retrieve-read" setting of the RAG paradigm.
 4.1 Experimental Setup
-We assess the robustness of RAG systems to spurious features by evaluating them on their most popular application—the Question Answering (QA) task, following the standard ”retrieve-read” setting of the RAG paradigm.
-We employ Contriever-msmarco as the default retriever and conduct our main experiments using the synthetic dataset generated by the SURE framework.
-Due to space limitations, additional implementation details can be found in Appendix F.
-4.2 Result Analysis
-In this section, we analyze our main results by category and offer insights from various perspectives.
-The results of Mistral-7B-Instruct and Llama-3.1-8B-Instruct are presented in Table 2 and Table 6, respectively.
-For Different Queries
-There is a significant difference in robustness rates between known and unknown queries when evaluated on noise documents (compare K-N and U-N), whereas no such gap is observed for golden documents (compare K-G and U-G).
-This phenomenon arises from the tug-of-war between an LLM’s internal prior and external evidence.
-When the retrieved content is noise documents, LLMs are likely to override it with their own correct prior knowledge (Known).
-However, the robustness to explicit noise may be affected by the spurious features in documents.
-In other words, implicit noise (spurious features) can coexist with explicit noise, highlighting the challenge of maintaining robustness to spurious features in complex noise scenarios.
-For Different Grounding Data
-Across all types of spurious features, the robustness rate on noise documents is consistently higher than that on golden documents, whether for known or unknown queries.
-This derives from the limitation of the RR metric we propose, which measures unrobustness by capturing changes in answer correctness rather than the inconsistency of responses.
-When tested on noise documents, regardless of which spurious features are injected, LLMs always generate incorrect responses, as noise documents lack ground truths.
-In this case, even though the responses change, the RR does not decrease since all responses remain incorrect.
-However, in real-world applications, our primary concern is whether the RAG system can consistently generate correct answer when faced with golden documents containing different spurious features.
-Therefore, we primarily focus on the RR results for the (K, G) and (U, G) subsets.
-For Different Metrics
-If the win rate surpasses the lose rate, it shows that more instances were corrected rather than misanswered after introducing perturbations.
-Based on the comparison of WR and LR in Table 2 and Table 6, we can conclude that Not every spurious feature is harmful and they can even be beneficial sometimes.
-For example, the WR of source perturbations is consistently higher than the LR on both models.
-This suggests that the performance of RAG can be enhanced by introducing beneficial spurious features (e.g., simply paraphrasing documents using LLMs).
+Datasets
+Through the steps of spurious features injection and causal features preservation, we derive the final dataset available for robustness evaluation: SURE_Wiki.
+The queries are drawn from the NQ-open dataset Lee et al.
+(2019), while our data source is English Wikipedia dump.
+Models
+We test two representative LLMs in our main experiments: Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct. Further implementation details are included in Appendix E.
+K-G K-N U-G U-N Total
+Style 7766 31152 2593 37692 79203
+Source 9249 32435 3228 39101 84013
+Logic 9724 35537 3587 41990 90838
+Format 11037 38018 4141 45518 98714
+Meta 11104 38018 4255 45420 98797
+Table 2: Statistics of the SURE_Wiki dataset for Mistral-7B-Instruct-v0.3.
+K-G denotes the instances composed of (Known query, Golden Document), while U-N refers to the instances consisting of (Unknown query, Noise Document).
+The values represents the number of instance pairs for each type of perturbations within the category of spurious features.
+4.2 Experimental Results
+To further analyze spurious features, we divide SURE_Wiki into four subsets based on the categories of queries and documents within each instance.
+A query is labeled as Known if it can be correctly answered in a closed-book setting; otherwise, it is labeled as Unknown.
+Documents are categorized as Golden or Noise depending on whether they contain ground truths.
+Notably, the distribution of the dataset is model-specific, as the classification of Known and Unknown queries is determined by the intrinsic knowledge of the target LLM.
+Table 2 presents dataset statistics for Mistral-7B-Instruct-v0.3, while the distribution for Llama-3.1-8B-Instruct is shown in Appendix D.
+For Different Queries and Grounding Data
+We report the results of Mistral-7B-Instruct and Llama-3.1-8B-Instruct in Table 1 and Table 9, respectively.
+For golden documents, the robustness rates of K-G and U-G are very similar for both Mistral and Llama, whereas their accuracy differ significantly.
+This suggests that, unlike robustness to explicit noise Wu et al.
+(2024b), robustness against spurious features is independent of the model’s internal prior knowledge.
+When tested on noise documents, the RR remains high across all spurious features, as LLMs consistently generate incorrect responses in the absence of ground truths.
+Therefore, we primarily focus on the RR results for the golden documents in the following experiments.
 For Different Perturbations
-We observe notable differences in the robustness rates among five types of spurious features.
-However, within each category, the robustness rates across different perturbations are relatively close.
-When further comparing perturbations within the same category, we find that while their RR values are comparable, their WR and LR can differ significantly, indicating LLMs’ preference for certain sub-perturbations.
-In summary, there are variations in robustness across different spurious features, and preference exists among the sub-perturbations of each spurious feature.
-4.3 SIG Benchmark & Further Analysis
+We observe notable differences in robustness rates across the five types of spurious features.
+However, within each category, the RR values for different perturbations are relatively similar.
+Hence, the robustness of spurious features can be estimated by averaging the RR values of their corresponding perturbations.
+When further comparing perturbations within the same category, we find that while their RR values are comparable, their WR and LR can differ significantly.
+If the WR exceeds the LR, more instances are corrected than misanswered after introducing perturbations.
+This suggests that not every spurious feature is harmful and they can even be beneficial sometimes.
+4.3 Further Analysis
 The raw synthetic dataset is not ideal for extensive evaluation due to its large size.
 Furthermore, the class imbalance result in unfair comparisons across different types of spurious features.
-To facilitate more efficient evaluation, we extract the most challenging data from the subsets of synthetic dataset to create a lightweight benchmark: SIG (Spurious features In Golden document).
-Specifically, we select instance pairs where both the Mistral and Llama models exhibit sensitivity and unrobustness in our main experiments.
-An equal number of samples (100) are chosen from the (K,G) and (U,G) subsets of each perturbations.
-Robustness Comparison of SOTA LLMs
-We evaluate a broader range of models on SIG benchmark, including GPT-4O, GPT-4O-mini, Mistral-Large-Instruct  [^5] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama-3.3-70B-Instruct, Deepseek-v3 (671B,MoE), and Qwen2.5-72B-Instruct.
+To facilitate more efficient evaluation, we extract the most challenging data from our synthetic datasets to create a lightweight benchmark: SIG_Wiki (Spurious features In Golden document from Wiki)  [^5] Specifically, we randomly select 100 instance pairs for each perturbation where both models lack robustness..
+Spurious Features in Different Model Families.
+To examine whether spurious features are merely artifacts of specific model choices, we evaluate a diverse set of SOTA LLMs on the SIG benchmark.
+The evaluated models include GPT-4O, GPT-4O-mini, Mistral-Large-Instruct  [^6] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama-3.3-70B-Instruct, Qwen2.5-72B-Instruct, and DeepSeek-V3 (671B,MoE), covering a wide range of model series and architectures.
 To better compare the robustness of different models, we average the RR of each perturbation within a category to derive the overall robustness for a specific type of spurious feature.
 The performance of six SOTA LLMs is then visualized using a radar chart, as shown in Figure 4.
-Notably, GPT-4O-mini achieved the best performance, even surpassing GPT-4O by a large margin on the format and meta types.
-Despite the impressive robustness of closed-source models, they may still exhibit sensitivity to certain specific perturbations.
-For instance, GPT-4o achieved only an 89% robustness rate on the datasource(twitter) perturbation.
-Figure 4: Comparison of Robustness rates across six SOTA LLMs.
-Scaling Analysis for Different Model Sizes
+Despite the impressive robustness of closed-source models, they still exhibit sensitivity to certain specific perturbations.
+These results demonstrate that spurious features are a widespread issue across different model families, sizes, and architectures (Dense VS.
+MoE).
+Figure 4: Robustness comparison of six SOTA LLMs.
+Spurious Features Across Retrieval Settings.
+To evaluate whether spurious features exist across different retrieval configurations (query sources, retrievers, and evidence sources), we additionally construct a benchmark, SIG_Trivial, following the SURE framework pipeline.
+This benchmark employs a production-level retriever (Bing Search) to retrieve documents from the open web, using queries from Trivial_QA Joshi et al.
+(2017).
+Further details are provided in Appendix F.
+Table 3 presents the average RR for each type of spurious feature.
+The results demonstrate that spurious features are a prevalent issue for RAG systems across different retrieval configurations.
+Style Source Logic Format Meta
+Mistral-7B-Instruct 88.0 94.0 94.5 94.0 99.0
+with LLM-as-Judge 90.5 91.5 92.0 93.8 96.0
+[1pt/1pt] Llama-3.1-8B-Inst. 87.5 93.5 93.0 90.8 97.0
+with LLM-as-Judge 85.0 92.0 91.0 90.8 93.3
+Table 3: Robustness evaluation of two models on SIG_Trivial, including results from LLM-as-Judge paradigm as a complementary assessment.
+Reliability of Robustness Evaluations.
+To validate the reliability of our string-based evaluation method, we introduce the LLM-as-Judge paradigm as a comparative assessment on SIG_Trivial benchmark.
+See more details in Appendix G.
+As illustrated in Table 3, our evaluation results similar to those of LLM-as-Judge, suggesting its effectiveness and efficiency.
+4.4 Mitigating Spurious Features
+Can Scaling up Model Size Solve the Problem?
 To investigate the impact of parameter scale on RAG robustness, we gradually increase the size of LLM-based readers (Qwen2.5 series, ranging from 0.5B to 72B) and evaluate their robustness across five types of spurious features.
@@
 However, when we further scale the model from 32B to 72B, the RR undergoes a significant decline (except for format and meta).
-This indicates that robustness issues related to spurious features cannot be resolved simply by increasing model size.
-Interestingly, for meta perturbations, while RALMs demonstrate strong robustness across all scales (even for the 0.5B model), their performance receives little to no benefit from scaling up.
-Figure 5: Scaling analysis of robustness to spurious features.
+These findings suggest that although scaling up model size can enhance robustness to some extent, it fails to fundamentally eliminate sensitivity to spurious features.
+Figure 5: Scaling analysis on Qwen2.5 series.
+Are Existing Prompting Methods Effective?
+We evaluate whether methods developed to improve the robustness of RALMs against explicit noise can generalize to spurious features.
+Previous work, such as Chain-of-Note (CON) Yu et al.
+(2023), aims to enhance robustness by generating thorough rationale before producing the answer.
+Moreover, recent breakthroughs in the reasoning capabilities of LLMs have significantly advanced the cutting edge of RAG.
+By integrating with reasoning models, RAG can overcome previous limitations and adapt to more complex scenarios Gao et al.
+(2025).
+Therefore, we test both CON and DeepSeel-R1 on our SIG benchmark (results shown in Table 4).
+Notably, the robustness rate of CON is even lower than the baseline (Qwen2.5-72B-Instruct) without applying CON.
+A similar phenomenon was observed in experiments with the reasoning model DeepSeek-R1 Guo et al.
+(2025), whose robustness was even worse than its base model, DeepSeek-V3.
+This indicates that the robustness against spurious features cannot be effectively improved through COT-style techniques.
+Style Source Logic Format Meta
+Qwen2.5-72B 78.5 76.0 88.6 92.5 95.0
++ Chain-of-Note 74.0 81.7 66.7 84.8 91.0
+[1pt/1pt] DeepSeek-V3 96.5 93.6 95.6 94.0 96.5
+DeepSeek-R1 84.5 87.3 83.3 87.0 87.5
+Table 4: Robustness evaluation of CoN and DeepSeek-R1. Values that show improvements over the baseline are marked in bold.
+Improving Robustness Using Synthetic Data.
+Using the data generated by our SURE framework, we introduce two training-based mitigation strategies, Supervised Finetuning (SFT) and Direct Preference Optimization (DPO), designed to enhance the robustness of RALMs against spurious features.
+Specifically, we select unrobust instances, each comprising a query, a correct answer, an incorrect answer, and two golden passages (original and perturbed documents).
+For SFT, we train the model to consistently produce the correct answer for each query, creating two training samples by pairing the answer separately with the original and perturbed golden passages.
+For DPO, we train the model to prefer the correct answer over the incorrect one, likewise constructing two samples per query using both golden passages.
+The goal of these strategies is to teach the model to reliably generate the correct answer, regardless of whether the input documents contain spurious features.
+We train our methods for two epochs on a dataset of over 30k samples, and evaluate their robustness rates on the SIG_Wiki and SIG_Trivial (out-of-domain) benchmarks.
+The backbone model is Llama-3.1-8B-Instruct.
+The details of the hyperparameter settings are presented in Appendix H.
+As shown in Table 5, both methods significantly enhance robustness.
+However, SFT performs better on in-domain data (i.e., SIG_Wiki), while DPO generalizes more effectively to out-of-domain data.
+These results demonstrate the effectiveness of our SURE framework for generating training data and provide strong baselines for future research.
+Style Source Logic Format Meta
+Llama3.1-8B (Wiki) 10.0 15.5 20.0 24.0 94.0
++ SFT 96.5 94.5 99.0 99.5 99.7
++ DPO 96.5 96.0 96.0 98.0 98.0
+[1pt/1pt] Llama3.1-8B (Trivial) 87.5 93.5 93.0 90.8 97.0
++ SFT 88.5 91.5 95.0 96.3 99.0
++ DPO 94.5 94.5 97.3 95.8 98.0
+Table 5: Robustness comparison of Llama-3.1-8B-Instruct trained with SFT and DPO.
+The upper section shows the results on SIG_Wiki, while the lower section presents the results on SIG_Trivial.
+The best score for each type is highlighted in bold.
 5 Conclusion
 In this work, we formally highlight the spurious features problem in RAG system.
-Through preliminary experiments, we provide statistical evidence to support the presence of spurious features in RALMs.
-We also propose a novel evaluation framework, SURE, to assess the robustness of LLMs against spurious features.
-This framework includes a comprehensive taxonomy of spurious features, carefully designed metrics for robustness evaluation, and a data synthesis pipeline.
-Extensive evaluation and further analysis offers valuable insights for future research.
-Impact Statement
-We introduce a new type of noise, spurious features, that can undermine the robustness and trustworthiness of contemporary RAG systems.
-With the increasing diversity of golden documents on the internet, this noise may become more prevalent.
-Furthermore, traditional methods for improving the robustness of RAG systems are ineffective for handling spurious features, as they are implicit and can coexist with golden documents.
-Therefore, harmful spurious features could undermines user trust and contaminates Internet data.
-Moving forward, we aim to explore methods for mitigating hallucinations and robustness issues caused by spurious features within grounding data.
-Appendix A Related Work
-A.1 Robustness Evaluation of Retrieval-Augmented Generation
-RAG systems comprise two core components: a retriever and an LLM-based reader.
-Augmenting LLMs with retrieved external knowledge has been proven to effectively reduce hallucinations (Shuster et al., 2021; Kang et al., 2023).
-However, the retrieved contexts inevitably contains noise in addition to desirable knowledge, which may mislead LLMs to produce an incorrect response (Bian et al., 2024; Feldman et al., 2024).
-Previous works have explored automated evaluation frameworks to assess the robustness of RAG systems in various settings.
-For instance, Chen et al.
-(2024a) benchmarked four fundamental capabilities required for RAG, including noise robustness, negative rejection, information integration and counterfactual robustness.
-Some studies have provided a detailed taxonomy of noise documents to further simulate the complexity of real-world scenarios and highlighted the potential positive effects of certain types of noise (Cuconasu et al., 2024; Wu et al., 2024a).
-There are also some recent works that propose using LLM-as-a-judge (Li et al., 2024a) to evaluate the RAG system (Wang et al., 2024b).
-While these studies have identified several explicit noises that affect the robustness of RAG systems, they overlook implicit noises.
-This type of noise, such as phrasing and formatting, is everywhere and unavoidable, as it coexists with the grounding data without altering its semantic information.
-In this work, we define these semantic-agnostic noises as spurious features and evaluate the robustness of RALMs to such noises.
-A.2 Prompt Sensitivity of LLMs
-Large Language Models take prompts as inputs and then generate response accordingly.
-Prompts are instructions provided to an LLM to perform specific tasks automatically and ensure desired qualities in the generated output.
-However, it is known that current LLMs are sensitive to the features of input prompts (Zhu et al., 2023).
-This sensitivity poses challenges for researchers attempting to evaluate the model’s performance accurately and precisely (Zhuo et al., 2024).
-Some existing works have investigated the impact of different prompt techniques on model performance, including chain-of-thought (Wei et al., 2022), in-context learning (Min et al., 2022), and role-play prompting (Kong et al., 2024).
-Beyond these causal features that significantly influence the meaning of prompts, other works have demonstrated that LLMs are highly sensitive to spurious features (Sclar et al., 2024), e.g, prompt formatting (He et al., 2024), language style (Li et al., 2023a), the order of options (Pezeshkpour & Hruschka, 2024).
-Currently, there is no statistical or empirical evidence to support the existence of spurious features in RALMs.
-To address this gap, we extend the definition of spurious features to RAG systems through statistical testing and empirical analysis.
-Appendix B Preliminary Experiment Results
-We introduce several semantic-agnostic features we measured in our preliminary experiments. These features include:
+To quantify the robustness of RALMs against spurious features, we propose a novel evaluation framework, SURE, which includes a comprehensive taxonomy, a data synthesis pipeline, and evaluation metrics.
+Furthermore, leveraging the synthetic data generated by SURE, we introduce two training-based approaches that effectively improve the robustness of RALMs.
+Overall, our framework enables the systematic evaluation and mitigation of spurious features, paving the way for future research.
+Limitations
+We strive to comprehensively cover all types of spurious features that may arise in RAG scenarios.
+However, some unidentifiable spurious features may fall outside the scope of our taxonomy and thus fail to be quantified using the proposed SURE framework.
+Ethical Considerations
+All datasets used in our work are publicly released under open licenses, and all models employed for generation are open-source and licensed for research use.
+We strictly follow the licensing terms and usage policies of these resources.
+Appendix
+Appendix A Preliminary
+In this section, we first define causal and spurious features in the context of RAG and then demonstrate the existence of spurious features statistically.
+A.1 Causal and Spurious Features in RAG
+In general, causal features are input features that have a direct causal effect on the output of predictive model Yu et al.
+(2020).
+Their relationship is rooted in causality, rather than mere statistical correlation.
+When it comes to Large Language Models, the meaning and intent of prompts serve as causal features that directly influence the models’ responses.
+In the context of RAG, causal features refer to the semantic information of grounding data.
+In contrast, spurious features are input features that co-occur with causal features and are erroneously captured by the model Neuhaus et al.
+(2023).
+These features exhibit a statistical correlation with the model’s output but lack a causal relationship.
+Recent research has shown that LLMs are sensitive to seemingly trivial features like prompt formatting, thereby extending the definition of spurious features to LLMs Sclar et al.
+(2024).
+Similarly, we define the semantic-agnostic features of the grounding data as spurious features in RAG systems.
+However, conclusions drawn from in-context learning scenarios (e.g., classification and multiple-choice tasks) may not applicable to RAG scenarios, which typically involve open-ended generation tasks.
+Therefore, we design a preliminary experiment to validate the presence of spurious features in RAG.
+A.2 Preliminary Experiment
+We aim to demonstrate the semantic-agnostic features within real documents are spurious features, i.e., to reveal their impact on the output of RAG systems.
+There are some challenges in revealing the influence of semantic-agnostic features.
+First, when retrieving from a single corpus, it is difficult to mine semantically equivalent counterparts that differ only in semantic-agnostic features.
+To mine appropriate documents, we introduce Contriever-msmarco, a traditional dense retriever, to recall 100 semantically similar candidates.
+To further eliminate the effect of causal features, documents without golden answers are filtered out, ensuring that the remaining documents have roughly consistent causal features.
+Still, the differences in spurious features among these candidate documents are often minor, and their impact on model responses cannot be effectively captured using binary evaluation methods that simply judge whether an answer is correct or incorrect.
+Thus, more fine-grained metrics are required to detect such nuanced performance changes.
+Inspired by the use of LLMs as supervision signals for document utility Izacard et al.
+(2023); Gan et al.
+(2024), we introduce the oracle score, which measures fine-grained performance through calculating the log probability of generating correct answers given a specific document.
+The oracle score is defined as follows:
+[…]
+where $x$ is the input prompt for RALMs, including the instruction $I$, grounding data $G$, and query $Q$; $y$ represents the ground truth answer; $\theta$ denotes the model parameters; and $T$ is the total length of the answer sequence  [^7] For cases with multiple answers, we compute the final score by averaging the corresponding oracle scores across all answers..
+For each query, we construct document pairs by selecting the first-ranked and last-ranked candidate documents based on their oracle scores.
+However, the presence of various semantic-agnostic features within each document pairs makes it challenging to isolate the impact of any individual features.
+To assess the influence of a given feature, we compare its distribution between document sets with first- and last-ranked oracle scores.
+A control group is constructed by randomly sampling two document sets.
+If the distributions differ significantly, it suggests that RALMs are sensitive to the feature.
+See Appendix A for implementation details of preliminary experiments.
+We test the following features: 1) Flesh Score, 2) Distinct-1, 3) Dependency Tree Depth, 4) PPL, and 5) Token Length.
+The results show that RALMs are sensitive to semantic-agnostic features.
+Nevertheless, it does not offer empirical evidence or quantitative analysis.
+Inspired by previous data synthesis studies Tan et al.
+(2024b), we use a data synthesis approach to better control feature variables and quantify the robustness of RALMs.
+A.3 Preliminary Experiment Results
+Using Contriever-msmarco, we recall 100 documents from the Wikipedia dump for each query in the NQ-open dataset.
+After filtering out documents that do not contain golden answers, we select the first-ranked and last-ranked documents based on their oracle scores for each query from the remaining documents, resulting in two sets of 2658 samples each.
+By comparing the differences in feature distributions between these two sets, we can assess whether RALMs exhibits sensitivity toward semantic-agnostic features.
+If these two sets do not belong to the same feature distribution, this can be attributed to the inherent bias of LLMs towards semantic-agnostic features.
+To confirm that this bias is not introduced by the dense retriever in first-stage retrieval, we establish a control group by randomly sampling two documents instead of selecting the first- and last-ranked documents.
+To evaluate whether the two distributions are same, we employ the Kolmogorov-Smirnov (K-S) test. The following semantic-agnostic features are measured in the experiments:
   • Flesch Score: A readability metric designed to evaluate text difficulty.
@@
 Otherwise, we fail to reject the null hypothesis, suggesting that there is no significant difference between the two distributions.
-We provide the K-S test results in Table 3 and Table 4, with the corresponding distribution visualizations plots shown in Figure 6 and Figure 7.
+The K-S statistic and P-value are presented in Table 6 and Table 7.
+Furthermore, we visualize the feature distributions for both the experimental and control groups in Figure 6.
+For all tested features in the experimental group, the K-S test rejects the null hypothesis, concluding that the distribution of the two sets are significantly different.
+In contrast, for the control group, the K-S test fails to reject the null hypothesis.
+The results for Llama-3.1-8B-Instruct are also provided in Figure 7.
+According to these results, we can conclude that RALMs exhibit bias toward spurious features in documents.
 Experimental Group Control Group
 K-S statistic P-value K-S statistic P-value
-[…]
-[…]
-[…]
-[…]
-[…]
+Flesch score 0.0677 $1.01\text{\times}{10}^{-5}$∗∗∗ 0.0301 0.1799
+Distinct-1 0.0756 $4.95\text{\times}{10}^{-7}$∗∗∗ 0.0203 0.6431
+DTD 0.0636 $4.29\text{\times}{10}^{-5}$∗∗∗ 0.0124 0.9866
+PPL 0.0722 $1.88\text{\times}{10}^{-6}$∗∗∗ 0.0162 0.8776
+Token Length 0.1708 $2.91\text{\times}{10}^{-34}$∗∗∗ 0.0256 0.3493
 Table 6: K-S test results for Mistral-7B-Instruct-v0.3 as the oracle retriever.
@@
 Flesch score 0.0305 0.1694 0.0173 0.8210
-[…]
+Distinct-1 0.0798 $8.94\text{\times}{10}^{-8}$∗∗∗ 0.0327 0.1159
 DTD 0.0474 0.0051∗∗ 0.0203 0.6431
 PPL 0.0538 0.0009∗∗∗ 0.0181 0.7791
-[…]
+Token Length 0.1275 $2.99\text{\times}{10}^{-19}$∗∗∗ 0.0188 0.7349
 Table 7: K-S test results for Llama-3.1-8B-Instruct as the oracle retriever.
@@
 Figure 7: Visualization of feature distributions for Llama-3.1-8B-Instruct
-Appendix C Implementation Details for Injecting Spurious Features
-We provide detailed prompts for LLM-based perturbations in Figure 8. For rule-based perturbations, placeholder template is presented in Figure 9.
+Appendix B Implementation Details for Injecting Spurious Features
+We provide detailed prompts for LLM-based perturbations in Figure 8. For rule-based perturbations, placeholder templates are presented in Figure 9.
 Style Perturbations [Simple] Please simplify the following text while preserving its original meaning.
@@
 Figure 9: Placeholder templates for rule-Based perturbations.
-Appendix D Implementation Details for Preserving Causal Features
-We employ a bidirectional entailment algorithm to ensure the semantic equivalence before and after introducing spurious features. The prompts for its core component, NLI model, are shown in Figure 10.
+Appendix C Implementation Details for Preserving Causal Features
+We employ a bidirectional entailment algorithm to ensure the semantic equivalence before and after introducing spurious features.
+The prompts for its core component, NLI model, are shown in Figure 10.
+Furthermore, we apply a simple string-matching strategy to preserve ground truths.
+Specifically, for Golden documents that originally contained the correct answers, we keep them only if they preserve the ground truths after perturbation.
+For Noise documents that initially lack the correct answers, we discard them if they unexpectedly acquire ground truths due to perturbations.
 Consider the two passages below.
@@
 Figure 10: Prompts for LLM-based NLI system.
-Appendix E Statistics of the Synthetic Dataset
+Appendix D Statistics of the Synthetic Dataset
 We present the dataset statistics for evaluating Llama-3.1-8B-Instruct in Table 8.
@@
 Meta 10563 35451 4796 47987 98797
-Table 5: Distribution of the synthetic dataset for Llama-3.1-8B-Instruct.
-Appendix F Experimental Setup Details
-Dataset
-We conduct our main experiments using the synthetic dataset generated by the SURE framework.
-The queries are sourced from the NQ-open dataset  [^6] https://huggingface.co/datasets/google-research-datasets/nq_open, and the documents are derive from the English Wikipedia dump.
-Models
-We test two representative LLMs in our main experiments: Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct.
+Table 8: Distribution of the SURE_Wiki dataset for Llama-3.1-8B-Instruct.
+Appendix E Experimental Setup Details
 Prompts
-The instruction $I$ in the RAG prompt $P=(I,G,Q)$, shown in Figure 11, is derived from Cuconasu et al. (2024), with slight modifications to better adapt to our setting.
+The instruction II in the RAG prompt $P=(I,G,Q)$, shown in Figure 11, is derived from Cuconasu et al. (2024), with slight modifications to better adapt to our setting.
 Implementation Details
-We follow the typical ”retrieve-read” setting of RAG paradigm.
+We follow the typical "retrieve-read" setting of RAG paradigm.
 For the retrieval module, we use Contriever-msmarco [^8] https://huggingface.co/facebook/contriever-msmarco, a BERT-based dense retriever, as the default retriever.
-It is finetuned on the MS MARCO dataset (Bajaj et al., 2016) after unsupervised pretraining via contrastive learning (Izacard et al., 2021).
-To optimize the efficiency of vector similarity searches, we employ the Faiss library (Douze et al., 2024).
+It is finetuned on the MS MARCO dataset Bajaj et al.
+(2016) after unsupervised pretraining via contrastive learning Izacard et al.
+(2021).
+To optimize the efficiency of vector similarity searches, we employ the Faiss library Douze et al.
+(2024).
 For the read module, we deploy LLMs on NVIDIA A100 GPUs and accelerate inference with vllm [^9] https://github.com/vllm-project/vllm.
-We follow (Li et al., 2023b; Zhang et al., 2024a, b) in using a greedy decoding strategy and set the temperature to 0.1 to ensure stable outputs and strong reproducibility.
+We set the temperature to 0.1 to ensure stable outputs and strong reproducibility.
 You are given a question and you MUST respond by EXTRACTING the answer (max 5 tokens) from the provided document. If the document does not contain the answer, respond with NO-RES.
-Figure 11: Instruction $I$ used for the QA task.
+Figure 11: Instruction II used for the QA task.
 Llama-3.1-8B-Instruct
-Spurious Features Perturbations Known-Golden Known-Noise Unknown-Golden Unknown-Noise Total
-LR RR WR LR RR WR LR RR WR LR RR WR LR RR WR
-Style Simple 7.79 83.04 9.18 1.70 95.80 2.50 8.43 82.88 8.69 0.27 99.45 0.28 1.80 95.96 2.24
-Complex 6.00 85.60 8.40 1.91 96.59 1.50 6.71 84.86 8.43 0.23 99.57 0.20 1.63 96.62 1.75
-Source LLM-Generated 5.89 86.43 7.69 1.43 96.83 1.74 6.20 85.71 8.09 0.22 99.56 0.22 1.51 96.60 1.89
-Self-Generated 6.55 85.01 8.44 1.55 96.37 2.09 6.52 86.36 7.12 0.19 99.57 0.24 1.62 96.32 2.06
-Logic Reverse 5.06 90.82 4.12 1.13 97.82 1.06 5.73 89.71 4.56 0.21 99.67 0.13 1.29 97.64 1.07
-Random 3.91 93.16 2.93 0.86 98.31 0.83 4.21 91.67 4.12 0.14 99.72 0.14 0.97 98.18 0.85
-LLM-Ranked 3.24 93.93 2.83 0.82 98.43 0.74 3.58 93.36 3.06 0.13 99.76 0.11 0.85 98.39 0.75
-Format JSON 7.01 88.25 4.74 1.70 97.25 1.05 5.92 89.63 4.45 0.25 99.61 0.14 1.76 97.08 1.16
-HTML 11.85 84.46 3.69 2.70 96.90 0.40 9.33 86.78 3.90 0.35 99.61 0.04 2.85 96.41 0.74
-YAML 5.26 89.94 4.80 1.26 97.41 1.33 4.79 90.80 4.41 0.17 99.67 0.16 1.32 97.40 1.28
-Markdown 2.32 92.23 5.45 0.60 96.89 2.51 2.34 93.46 4.19 0.07 99.61 0.32 0.61 97.55 1.84
-Metadata Timestamp (pre) 2.08 95.81 2.11 0.28 99.42 0.29 2.54 95.56 1.90 0.02 99.95 0.03 0.46 99.10 0.44
-Timestamp (post) 2.04 95.86 2.10 0.25 99.43 0.32 2.81 95.56 1.63 0.02 99.95 0.03 0.46 99.11 0.43
-Datasource (wiki) 2.11 93.45 4.44 0.23 98.96 0.81 3.25 92.47 4.27 0.03 99.86 0.11 0.48 98.50 1.03
-Datasource (twitter) 2.27 94.11 3.62 0.31 99.25 0.43 2.77 93.97 3.25 0.02 99.91 0.07 0.50 98.77 0.73
+Taxonomy Perturbations Known-Golden Known-Noise Unknown-Golden U-N
+LR $R$ WR Org Acc LR $R$ WR Org Acc LR $R$ WR Org Acc $R$
+Style Simple 7.79 83.04 9.18 66.03 67.42 1.70 95.80 2.50 4.12 4.92 8.43 82.88 8.69 51.42 51.68 99.45
+Complex 6.00 85.60 8.40 68.43 1.91 96.59 1.50 3.71 6.71 84.86 8.43 53.13 99.57
+[1pt/1pt] Source LLM-Generated 5.89 86.43 7.69 65.62 67.43 1.43 96.83 1.74 4.13 4.45 6.20 85.71 8.09 49.15 51.04 99.56
+Self-Generated 6.55 85.01 8.44 67.52 1.55 96.37 2.09 4.67 6.52 86.36 7.12 49.74 99.57
+[1pt/1pt] Logic Reverse 5.06 90.82 4.12 62.95 62.01 1.13 97.82 1.06 4.43 4.36 5.73 89.71 4.56 45.84 44.66 99.67
+Random 3.91 93.16 2.93 61.97 0.86 98.31 0.83 4.40 4.21 91.67 4.12 45.74 99.72
+LLM-Ranked 3.24 93.93 2.83 62.54 0.82 98.43 0.74 4.36 3.58 93.36 3.06 45.32 99.76
+[1pt/1pt] Format JSON 7.01 88.25 4.74 63.91 61.64 1.70 97.25 1.05 3.87 3.21 5.92 89.63 4.45 49.35 47.88 99.61
+HTML 11.85 84.46 3.69 55.75 2.70 96.90 0.40 1.56 9.33 86.78 3.90 43.92 99.61
+YAML 5.26 89.94 4.80 63.45 1.26 97.41 1.33 3.94 4.79 90.80 4.41 48.97 99.67
+Markdown 2.32 92.23 5.45 67.04 0.60 96.89 2.51 5.77 2.34 93.46 4.19 51.20 99.61
+[1pt/1pt] Metadata Timestamp (pre) 2.08 95.81 2.11 55.77 55.80 0.28 99.42 0.29 1.58 1.59 2.54 95.56 1.90 43.31 42.66 99.95
+Timestamp (post) 2.04 95.86 2.10 55.84 0.25 99.43 0.32 1.64 2.81 95.56 1.63 42.12 99.95
+Datasource (wiki) 2.11 93.45 4.44 58.10 0.23 98.96 0.81 2.17 3.25 92.47 4.27 44.33 99.86
+Datasource (twitter) 2.27 94.11 3.62 57.11 0.31 99.25 0.43 1.70 2.77 93.97 3.25 43.79 99.91
 Table 9: Robustness evaluation results of Llama-3.1-8B-Instruct on the synthetic dataset. We use Bold to mark the WR values that are higher than the LR, suggesting that the perturbation is beneficial.
+Appendix F Details of SIG_Trivial Benchmark
+We additionally construct a new benchmark, SIG_Trivial, following the SURE framework pipeline.
+This new benchmark uses a production-level retriever (Bing Search) to retrieve documents from the open web, with queries sources, retrieval algorithms, and data sources all differing from those used in our original SIG dataset.
+Specifically, we sample 1000 queries from TriviaQA and then sample 3 web documents (longer length than passages from wikepdeia) with golden answers from Bing Search for each query to serve as the grounding data.
+For each subcategory, 100 instances are randomly sampled to construct the new benchmark.
+Appendix G Implementation of LLM-as-Judge
+Specifically, we provide two demonstrations and prompt the LLM (Llama-3.1-70B-Instruct) to act as an evaluator.
+Before reaching a final judgment, the model is instructed to reason step by step using the chain-of-thought (COT) approach, encouraging a more thorough consideration of answer quality.
+Appendix H Training Setups
+Table 10 presents the detailed hyper-parameters for our training. We conduct our experiments on a server equipped with eight NVIDIA A100-SXM4-80GB GPUs.
+Hyper-parameter Value
+Learning Rate $2\times 10^{-5}$
+Number of Epochs 2
+Number of Devices 4
+Per-device Batch Size 4
+Gradient Accumulation Steps 2
+Effective Batch Size 32
+Optimizer AdamW
+Learning Rate Scheduler cosine
+Warmup Steps 100
+Max Sequence Length 4096
+Table 10: The hyper-parameters used for full parameter fine-tuning of models.